### v2

# references.bib

Abstract.
We classify $n$-representation infinite algebras $\Lambda$ of type $\tilde{A}$. This type is defined by requiring that $\Lambda$ has higher preprojective algebra $\Pi_{n+1}(\Lambda)\simeq k[x_{1},\ldots,x_{n+1}]\ast G$, where $G\leq\operatorname{SL}_{n+1}(k)$ is finite abelian. For the classification, we group these algebras according to a more refined type, and give a combinatorial characterisation of these types. This is based on so-called height functions, which generalise the height function of a perfect matching in a dimer model. In terms of toric geometry and McKay correspondence, the types form the lattice simplex of junior elements of $G$. We show that all algebras of the same type are related by iterated $n$-APR tilting, and hence are derived equivalent. By disallowing certain tilts, we turn this set into a finite distributive lattice, and we construct its maximal and minimal elements.
1991 Mathematics Subject Classification:
16G20, 16S35, 16E65, 14E16
Introduction
The famous McKay correspondence, named after John McKay who first discovered it [MR0604577], gives a bijection between finite subgroups of $\operatorname{SL}_{2}(\mathbb{C})$ and simply laced Dynkin diagrams. It can be realised geometrically by studying the associated quotient singularities $\mathbb{C}^{2}/G$ and their resolution graphs, and relating them to the McKay quivers $Q(G)$ of the involved groups. The McKay quiver $Q(G)$ turns out to be a doubled version of the corresponding extended Dynkin diagram.
Another interpretation involves the skew-group algebra $\mathbb{C}[x,y]\ast G$, which is Morita equivalent to a quotient of the path algebra $kQ(G)$ by commutativity relations $I$. This quotient $kQ(G)/I$ is the preprojective algebra of an acyclic orientation of the underlying extended Dynkin diagram. In fact, different acyclic orientations yield the same preprojective algebra, so via the preprojective algebra every basic connected tame hereditary algebra appears in the correspondence.
It is of great interest to generalise the correspondence, or at least parts of it, to higher dimensions. Geometrically, one can consider finite subgroups of $\operatorname{SL}_{n+1}(\mathbb{C})$. On the representation-theoretic side, a generalisation of the situation is provided by Iyama’s higher Auslander-Reiten theory (AR-theory).
The higher versions of hereditary representation infinite algebras are called $n$-representation infinite algebras, which were introduced by Herschend, Iyama and Oppermann in [HIO]. In the same paper, the authors also introduced $n$-hereditary algebras, and proved in a dichotomy-style theorem that this class of algebras naturally splits into $n$-representation infinite algebras and $n$-representation finite algebras.
The $n$-representation finite algebras and the related $n$-cluster tilting subcategories have been studied extensively since their introduction [Herschend_Iyama_2011, IyamaOppermann, IyamaOppermannStable, DarpöIyama]. The infinite side has received less attention, and fewer examples are known, partly due to the difficulty of checking directly whether a given algebra satisfies the definition of being $n$-representation infinite. There are some general constructions like tensor products [HIO] and $n$-Auslander-Platzeck-Reiten ($n$-APR) tilting [MizunoYamaura] which produce $n$-representation infinite algebras directly, and they may be obtained by finding suitable tilting objects in derived categories of varieties [BuchweitzHille].
The main source, however, is the class of $(n+1)$-Calabi-Yau algebras, taking a detour via the higher preprojective algebras: In detail, to an $n$-hereditary algebra $\Lambda$, Oppermann and Iyama [IyamaOppermannStable] associate the higher preprojective algebra as the tensor algebra
$\Pi_{n+1}(\Lambda)=T_{\Lambda}\operatorname{Ext}^{n+1}(D(\Lambda),\Lambda),$
which naturally comes with a grading induced from tensor degrees. Then one obtains $\Lambda$ as the degree 00 part of $\Pi_{n+1}(\Lambda)$. It was shown [KellerVandenBergh, HIO] that if $\Lambda$ is $n$-representation infinite, then $\Pi_{n+1}(\Lambda)$ is locally finite dimensional and bimodule $(n+1)$-Calabi-Yau of Gorenstein parameter $1$. Furthermore, these properties characterise all $n$-representation infinite algebras. Thus, one may take an ungraded $(n+1)$-Calabi-Yau algebra and produce a grading of Gorenstein parameter $1$ to construct an $n$-representation infinite algebra. This leads to the following general question [Thibault].
Question.
Let $\Gamma$ be bimodule $(n+1)$-Calabi-Yau. When does there exist a grading making $\Gamma$ into an $(n+1)$-preprojective algebra?
When the Calabi-Yau algebra $\Gamma$ is basic, the gradings we refer to are called cuts. Performing an $n$-APR tilt then amounts to a change of grading called cut mutation [IyamaOppermann]. We follow this strategy of computing preprojective algebras, and use it to give a description of all $n$-representation infinite algebras of type $\tilde{A}$. This type is defined by analogy with the classical McKay correspondence. A $1$-representation infinite algebra $\Lambda$ is nothing but a hereditary representation infinite algebra, and type $\tilde{A}$ means that $\Lambda\simeq kQ$ where $Q$ is an acyclic orientation of an extended Dynkin diagram of type $\tilde{A}$. In this case, the preprojective algebra $\Pi_{2}(kQ)$ is isomorphic to a skew-group algebra $k[x_{1},x_{2}]\ast G$, where $G\leq\operatorname{SL}_{2}(k)$ is a finite abelian group acting on the variables of $k[x_{1},x_{2}]$. The higher case is defined by requiring that $\Lambda$ is $n$-representation infinite and that $\Pi_{n+1}(\Lambda)\simeq R\ast G$, where $G\leq\operatorname{SL}_{n+1}(k)$ is a finite abelian group acting on $R=k[x_{1},\ldots,x_{n+1}]$.
The first subtlety that arises with this definition is that not every such skew-group algebra $R\ast G$ will appear as a higher preprojective algebra [Thibault]. Additionally, it is a priori not obvious how to decide whether a given $R\ast G$ does appear, and if so how to find the correct grading that will make it into a higher preprojective algebra. Furthermore, there may be several different gradings that turn $R\ast G$ into a higher preprojective algebra. We investigated this question for the case $n=2$ in [DramburgGasanova] and gave a classification for type $\tilde{A}$. The case $n=2$ has also been studied from a more geometric perspective by Nakajima using dimer models in [Nakajima].
We exploit the same ideas we used for $n=2$ in [DramburgGasanova], and show that they generalise to arbitrary $n$. More precisely, we assign a more refined type to the cuts of interest. We then show that these types naturally are lattice vectors in some simplex $P$, see 3.5. Deciding whether a given algebra $R\ast G$ has a preprojective cut is then equivalent to deciding whether $P$ has an internal lattice point, and the negative case splits naturally into some lower dimensional families and finitely many exceptions, see 3.11.
The analogy with McKay correspondence then continues as follows. Since $G$ is abelian, the variety $X=\operatorname{Spec}(R^{G})$ associated to the invariant ring is toric, hence defined by a lattice cone $C$. The simplex $P$, which we obtain in a purely combinatorial way, turns out to be isomorphic to the simplex one obtains by intersecting $C$ with a suitable hyperplane at “height $1$”, see 4.5. In standard toric terms, this can be seen as a consequence of a cut, that is a grading, on $R\ast G$ defining a certain $1$-parameter subgroup of $X$. Using the language of Ito and Reid from [ItoReid], we can see that our types correspond to junior elements in $G$, which are known to form the same simplex we have been considering, see 4.7. This completes the pictures, since our initial construction, following Amiot, Iyama and Reiten [AIR], is based on finding certain junior elements. The geometric perspective shows that the decision problem of whether $R\ast G$ has such a higher preprojective grading is equivalent to deciding whether $G$ contains certain nontrivial junior elements, which in turn correspond to exceptional crepant divisors in a relative minimal model of $X$. While this is of course computable for a given $G$, we are not aware of a simple criterion similar to [Thibault, Theorem 5.15] that allows us to decide the existence of such a junior element without computing all elements of $G$, or the simplex $P$, or some geometric data. We furthermore note that the quivers, and in particular the degree 00 parts of our cuts, have been recently considered by Yamagishi in [YamagishiIAbelian]. There, Yamagishi works with $G$-constellations and their moduli spaces for a finite abelian group $G\leq\operatorname{SL}_{n+1}(k)$. The motivation is purely geometric, and we are not aware of an a priori reason for why quivers of $n$-representation infinite algebras should appear.
After fixing an algebra $R\ast G$ and a type, we investigate how to obtain all the cuts of this type. We show that cut mutation is transitive on the set of cuts of a given type, see 5.12, which was already known for $n=2$ from the dimer perspective [Nakajima] or from combinatorial considerations [DramburgGasanova]. In this way, we answer a question of Amiot, Iyama and Reiten [AIR] by showing that all cuts are mutation equivalent to cuts arising from cyclic groups.
This also leads to an algorithmic classification of $n$-representation infinite algebras of type $\tilde{A}$: For each abelian group $G$ of order $m$, we can go through all embeddings of $G$ into $\operatorname{SL}_{n+1}(k)$ by considering generating sets of $\hat{G}$ of size $n$. In the next step, for a fixed embedding we can calculate the simplex $P$ of types of cuts. Then for each type, we can construct a cut. Finally, we can apply mutations until we found all cuts of the same type. In this way, we can list all type $\tilde{A}$ algebras with $m$ simple modules.
Our transitivity result is part of a broader observation. The set of cuts of a fixed type can be given the structure of a finite distributive lattice, whose cover relations are given by certain mutations. More precisely, we fix a vertex in the quiver $Q$ of $R\ast G$ at which we do not allow mutations, and the remaining mutations give the cover relations, see 5.7 and 5.11. One way to view this lattice is by considering the height functions on $Q$, which we define in Section 2. These functions generalise the $n=2$ theory of dimer models, and they turn out to capture our combinatorics in any dimension. Since $Q$ is finite, the height functions are determined by finitely many values, so they can be seen as vectors in some $\mathbb{Z}^{m}$, and the lattice of cuts of a fixed type translates via height functions to a sublattice of $(\mathbb{Z}^{m},\leq)$. A more thorough investigation of these lattices seems warranted, and we initiate this by constructing their minimal and maximal elements.
We may view $Q$ as being embedded on the $n$-torus, and the height functions as a replacement for perfect matchings of the dimer model. This idea was previously considered from a statistical point of view in [LammersDimer] for bounded regions. It would be interesting to generalise our observations to quivers of other $(n+1)$-Calabi-Yau algebras.
Outline
We begin by recalling some facts about $n$-representation infinite algebras, their preprojective algebras, quiver descriptions, higher preprojective gradings and skew-group algebras in Section 1. The first technical tool we need is called the height function, which is the main focus in Section 2, and it is used to prove one implication of our first main theorem 3.5 in 2.13. In Section 3, we then prove the other implication via an explicit construction in 3.4, and comment on the problem of deciding the existence of cuts in Section 3.2. After that, we reinterpret 3.5 and give explicit connections to toric geometry in Section 4. Finally, in Section 5, we consider cut mutation. In 5.8, we show that cuts of the same type form a finite distributive lattice, and obtain in 5.12 that they are related by a sequence of mutations. We construct the maximal element of each lattice in 5.17.
Notation
We use the following notation and conventions throughout.
  1. (1)
    Unless specified otherwise, we denote by $k$ an algebraically closed field of characteristic 00.
  2. (2)
    All undecorated tensors $\otimes$ are taken over the field $k$.
  3. (3)
    We denote by $\operatorname{diag}(x_{1},\ldots,x_{n})$ the diagonal $(n\times n)$-matrix with diagonal entries $x_{1},\ldots,x_{n}$.
  4. (4)
    We write $\frac{1}{m}(e_{1},\ldots,e_{n})$ for the diagonal matrix $\operatorname{diag}(\xi^{e_{1}},\ldots,\xi^{e_{n}})$ of order $m$, where $\xi$ is a primitive $m$-th root of unity which will be specified when necessary.
  5. (5)
    A quiver $Q$ consists of vertices $Q_{0}$, arrows $Q_{1}$ and the source and target maps $s,t\colon Q_{1}\to Q_{0}$.
  6. (6)
    For $a\in\mathbb{Z}$ and $n\in\mathbb{N}$, we write $a\bmod n$ for the smallest non-negative representative of the residue class $a+n\mathbb{Z}$.
  7. (7)
    For $a,b\in\mathbb{Z}$ and $n\in\mathbb{N}$ we write $a\equiv_{n}b$ if $a+n\mathbb{Z}=b+n\mathbb{Z}$. We drop the subscript $n$ when no confusion is possible.
1. Setup
In this section, we collect the necessary definitions and tools we will use throughout. We work over an algebraically closed field $k$ of characteristic 00.
1.1. Higher representation-infinite algebras and their preprojective algebras
Let $\Lambda$ be a finite dimensional $k$-algebra of finite global dimension $n\geq 1$. In Iyama’s higher Auslander-Reiten theory, the following functors play an important role:
$\nu=D\operatorname{\mathbb{R}\kern-1.99997pt\operatorname{Hom}}_{\Lambda}(\--,\Lambda)\colon\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda)\to\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda)$
is the derived Nakayama functor on the bounded derived category of finitely generated $\Lambda$-modules, and its quasi-inverse is
$\nu^{-1}=\operatorname{\mathbb{R}\kern-1.99997pt\operatorname{Hom}}_{\Lambda^{\mathrm{op}}}(D(\--),\Lambda)\colon\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda)\to\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda).$
The derived higher Auslander-Reiten translation is the autoequivalence
$\nu_{n}:=\nu\circ[-n]\colon\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda)\to\operatorname{\mathrm{D}^{\mathrm{b}}}(\operatorname{mod}\Lambda).$
Definition 1.1.
[HIO, Definition 2.7] The algebra $\Lambda$ is called $n$-representation infinite if for any projective module $P$ in $\operatorname{mod}\Lambda$ and integer $i\geq 0$ we have
$\nu_{n}^{-i}P\in\operatorname{mod}\Lambda.$
For a given $\Lambda$, this condition may be difficult to check since it requires knowledge of the infinitely many translates $\nu_{n}^{-i}(\Lambda)$. However, the higher preprojective algebra allows a different characterisation.
Definition 1.2.
[IyamaOppermannStable, Definition 2.11] Let $\Lambda$ be of global dimension at most $n$. The $(n+1)$-preprojective algebra of $\Lambda$ is
$\Pi_{n+1}(\Lambda)=T_{\Lambda}\operatorname{Ext}^{n}_{\Lambda}(D(\Lambda),\Lambda).$
We will refer to such an algebra $\Pi_{n+1}(\Lambda)$ simply as preprojective. The preprojective algebra comes naturally equipped with a grading, given by tensor degree, and we recover the original algebra as the degree 00 part, i.e. we have $\Lambda=\Pi_{n+1}(\Lambda)_{0}$. The properties of this grading, together with being Calabi-Yau, essentially determine such preprojective algebras.
Definition 1.3.
[AIR, Definition 3.1] Let $\Gamma=\bigoplus_{i\geq 0}\Gamma_{i}$ be a positively graded $k$-algebra. We call $\Gamma$ a bimodule $(n+1)$-Calabi-Yau algebra of Gorenstein parameter $a$ if there exists a bounded graded projective $\Gamma$-bimodule resolution $P_{\bullet}$ of $\Gamma$ and an isomorphism of complexes of graded $\Gamma$-bimodules
$P_{\bullet}\simeq\operatorname{Hom}_{\Gamma^{\mathrm{e}}}(P_{\bullet},\Gamma^{\mathrm{e}})[n+1](-a).$
Note that if $\Gamma$ is bimodule $(n+1)$-Calabi-Yau of Gorenstein parameter $a$, then the ungraded version of $\Gamma$ is simply $(n+1)$-Calabi-Yau. Conversely, different gradings may be introduced on $\Gamma$, so that the projective bimodule resolution becomes graded, and we can study the resulting Gorenstein parameter. The relevance of the Gorenstein parameter comes from the following theorem.
Theorem 1.4.
[HIO, Theorem 4.35] There is a bijection between $n$-representation infinite algebras $\Lambda$ and bimodule $(n+1)$-Calabi-Yau algebras $\Gamma$ of Gorenstein parameter 1 with $\dim_{k}\Gamma_{i}<\infty$ for all $i\in\mathbb{N}$, both sides taken up to isomorphism. The bijection is given by
$\Lambda\mapsto\Pi_{n+1}(\Lambda)\quad\mathrm{and}\quad\Gamma\mapsto\Gamma_{0}.$
In particular, this means that being higher preprojective is the additional structure of a specific grading on a Calabi-Yau algebra. Just like the term graded algebra means an algebra $\Gamma$ with a chosen grading, we therefore mean by preprojective algebra an algebra $\Gamma$ with a chosen grading such that $\Gamma=\Pi_{n+1}(\Gamma_{0})$. We can therefore phrase the general motivating question as follows.
Question 1.5.
Let $\Gamma$ be bimodule $(n+1)$-Calabi-Yau. Is $\Gamma$ higher preprojective? How many different gradings exist that make $\Gamma$ higher preprojective?
We restrict our interest to the so-called type $\tilde{A}$ in this article. The type $\tilde{A}$ is also defined in terms of the preprojective algebra. For this definition, we briefly recall the construction of a skew-group algebra. Given a $k$-algebra $A$ and a group $G$ acting on $A$ via automorphisms, the skew-group algebra is the tensor product
$A\otimes kG$
with multiplication given by
$(a\otimes g)\cdot(b\otimes h)=(ag(b)\otimes gh).$
In particular, given a group $G\leq\operatorname{SL}_{n+1}(k)$, we have that $G$ acts by left-multiplication on $V=\langle x_{1},x_{2},\ldots,x_{n+1}\rangle$, and hence on the polynomial ring $R=k[V]$.
Definition 1.6.
An $n$-representation infinite algebra $\Lambda$ has type $\tilde{A}$ if $\Pi_{n+1}(\Lambda)\simeq k[x_{1},\ldots,x_{n+1}]\ast G$ for a finite abelian group $G\leq\operatorname{SL}_{n+1}(k)$.
For the purposes of classifying the $n$-representation infinite algebras we may therefore first compute $R\ast G$, and then compute all possible higher preprojective gradings. However, there are many examples of finite abelian groups $G\leq\operatorname{SL}_{n+1}(k)$ such that $R\ast G$ can not be endowed with this additional structure, and we are not aware of an easy criterion to decide when this is possible. We note an important result in this direction, proven by Thibault also for non-abelian groups.
Theorem 1.7.
[Thibault, Theorem 5.15] Suppose that the embedding of a finite subgroup $G\leq\operatorname{SL}_{n+1}(k)$ factors through some embedding of $\operatorname{SL}_{n_{1}}(k)\times\operatorname{SL}_{n_{2}}(k)$ into $\operatorname{SL}_{n+1}(k)$ for $n_{1}+n_{2}=n+1$. Then $R\ast G$ can not be endowed with the grading of a $(n+1)$-preprojective algebra.
However even for abelian groups, there are examples showing that this criterion is not sufficient, see [DramburgGasanova, Example 3.15, Example 6.5]. To give some constructions of preprojective structures, we need an explicit quiver description of $R\ast G$, and we need to describe how it interacts with preprojective gradings. Before we do so, we need the following simplification of the setup.
1.2. Cuts
Note that for an abelian group $G$, the algebra $R\ast G$ is basic and can therefore be written as the quotient of a quiver algebra, i.e. we have $R\ast G\simeq kQ/I$. To construct a grading, it then often suffices to grade the quiver $Q$, i.e. to place certain arrows in $Q$ in degree $1$ so that $kQ/I$ becomes graded. This method has been employed e.g. in [Giovannini, DramburgGasanova, HIO]. However, it is a priori not clear that every possible higher preprojective grading will arise in this way. We therefore make the following definition.
Definition 1.8.
Let $R\ast G\simeq kQ/I$. A higher preprojective grading on $R\ast G$ is called a cut if every arrow in $Q$ is homogeneous of degree 00 or $1$. For such a cut, we let $C=\{a\in Q_{1}\mid\deg(a)=1\}$ and refer to the set $C$ as a cut as well.
Note that not every higher preprojective grading is a cut. For a detailed discussion of this problem, we refer the reader to [DramburgGasanova, Section 3.1]. However, it follows from [DramburgSandoy, Theorem 5.13] that every higher preprojective grading on $R\ast G$ can be mapped to a cut by an automorphism. In particular, this means that every $n$-representation infinite algebra $\Lambda$ with $\Pi_{n+1}(\Lambda)\simeq R\ast G\simeq kQ/I$ can be realised as the degree 00 part of a cut for a fixed $Q$. Taking the preprojective degree 00 part of an algebra with a cut amounts to removing, or cutting, all arrows in $C$ from the quiver $Q$ and removing all relations from $I$ which contain arrows in $C$. In order to refer to this procedure, the following notation will be used throughout.
Definition 1.9.
Let $C\subseteq Q_{1}$ be a cut. Then we refer to the cut quiver $Q_{C}=(Q_{0},Q_{1}-C)$ and the cut quiver algebra $(kQ/I)_{C}=\langle k(Q_{C})\rangle\leq kQ/I$.
1.3. Quivers and universal covers
In this subsection, we give a detailed summary of the quiver description of $R\ast G$ for abelian $G$. This is based on [BSW] and [HIO]. See also [DramburgGasanova, Section 7.1] for a similar summary. We recall that $G$ acts on $V=\langle x_{1},\ldots,x_{n+1}\rangle$ via the embedding of $G$ in $\operatorname{SL}_{n+1}(k)$. In particular, we view $V$ as a representation of $G$. Since $G$ is abelian, we can diagonalise this representation. That means decompose $V=\rho_{1}\oplus\cdots\oplus\rho_{n+1}$ into 1-dimensional representations so that the subspace $\langle x_{i}\rangle$ affords the representation $\rho_{i}$. As proven in [BSW], the quiver for $R\ast G$ is the McKay quiver of $G$. That means $R\ast G\simeq kQ/I$, where $Q=(Q_{0},Q_{1})$ has vertices $Q_{0}=\operatorname{Irr}(G)$ given by the irreducible representations of $G$. For $\chi_{1},\chi_{2}\in\operatorname{Irr}(G)$, the arrows $\chi_{1}\to\chi_{2}$ are a basis of $\operatorname{Hom}_{G}(\chi_{1},\chi_{2}\otimes V)$. Since $V$ is a faithful representation, it follows that $\{\rho_{1},\ldots,\rho_{n}\}$ is a generating set for the dual group $\hat{G}=\operatorname{Hom}(G,\mathbb{C}^{\ast})$ [DramburgGasanova, Lemma 4.2]. We arrive at the following observation.
Remark 1.10.
Let $G$ be abelian. The quiver $Q$ for $R\ast G$ is the directed Cayley graph of $\hat{G}$ with respect to the generating set determined by $V$. In particular, every vertex of $Q$ has $n+1$ outgoing and $n+1$ incoming arrows.
The fact that $Q$ is a Cayley graph also suggests that we associate a type to each arrow, corresponding to the $\rho_{i}$ it is associated to.
Definition 1.11.
Let $Q$ be the quiver of $R\ast G$, where $G$ is abelian. Every arrow $a\colon\chi_{1}\to\chi_{2}$ corresponds to an isomorphism $\chi_{1}\simeq\chi_{2}\otimes\rho_{i}$ for a unique $11\leq i\leq n+$ 1, and we call $\theta(a)=i$ the type of the arrow.
We recall from [BSW] that the relations $I$ for $R\ast G\simeq kQ/I$ are generated by commutator relations.
Remark 1.12.
The isomorphism $R\ast G\simeq kQ/I$ takes the element $x_{i}\otimes 1\in R\ast G$ to the sum of all arrows of type $i$. The commutativity relations $x_{i}x_{j}=x_{j}x_{i}$ in $R$ give rise to the defining relations of $kQ/I$ as follows. Let $a_{i}\colon x\to y$ and $a_{j}\colon y\to z$ be of types $\theta(a_{i})=i\neq j=\theta(a_{j})$. Then there exist arrows $b_{j}\colon x\to y^{\prime}$ and $b_{i}\colon y^{\prime}\to z$ of types $\theta(b_{j})=j$ and $\theta(b_{i})=i$. In $kQ/I$, we have
$a_{i}a_{j}=b_{j}b_{i},$
and $I$ is generated by all relations of this form.
Next, we describe a universal cover of $Q$ as detailed in [HIO, Section 5]. We denote the standard basis of $\mathbb{R}^{n}$ by $\{\alpha_{1},\ldots,\alpha_{n}\}$ and fix $\alpha_{n+1}=-\sum_{i=1}^{n}\alpha_{i}$. We consider the $n$-dimensional lattice
$L_{0}=\langle\alpha_{i}\mid 1\leq i\leq n\rangle$
and the corresponding infinite quiver
$\hat{Q}=(L_{0},\{x\to x+\alpha_{i}\mid 1\leq i\leq{n+1}\}).$
To cover $Q$, we consider the map
$L_{0}\to\hat{G},\alpha_{i}\mapsto\rho_{i},$
where $V=\rho_{1}\oplus\cdots\oplus\rho_{n+1}$ is the decomposition of the representation $V$. This is a group homomorphism, which on the level of quivers induces a surjection
$\hat{Q}\to Q,$
such that the arrows $x\to x+\alpha_{i}$ in $\hat{Q}$ map to the arrows of type $i$ in $Q$. Completing this into a short exact sequence of abelian groups
$0\to L_{1}\to L_{0}\to\hat{G}\to 0,$
we see that the kernel $L_{1}$ is itself an $n$-dimensional lattice, so we may think of the quiver $Q$ being embedded on an $n$-torus, while $\hat{Q}$ is the universal cover of this quiver in $\mathbb{R}^{n}$. The following remark shows that we can fix $L_{0}$ and then use the data of $L_{1}$ in place of considering finite subgroups of $\operatorname{SL}_{n+1}(k)$.
Remark 1.13.
The datum of a faithful $n$-dimensional representation $\rho$ of an abelian group $G$ is equivalent to the datum of a generating set $\{\rho_{1},\ldots,\rho_{n}\}$ of $\hat{G}$. This way, we obtain an equivalence between abelian subgroups $G\leq\operatorname{SL}_{n+1}(k)$ and cofinite sublattices $L_{1}\leq L_{0}$.
We now use the cover $\hat{Q}$ to fix important terminology.
Definition 1.14.
We keep the setup for $\hat{Q}$.
  1. (1)
    We say that an arrow $x\to x+\alpha_{i}$ has type $i$.
  2. (2)
    A cycle of length $n+1$ in $\hat{Q}$ consisting of arrows of $n+1$ distinct types is called an elementary cycle.
  3. (3)
    We call a set of arrows $\hat{C}\subseteq\hat{Q}_{1}$ a cut if every elementary cycle passes through exactly one arrow in $\hat{C}$.
  4. (4)
    The same terminology applies to arrows and cycles in $Q$.
  5. (5)
    We call a cut in $\hat{Q}$ periodic with respect to $L_{1}$ if it is invariant under $L_{1}$-translation, that is, $x\to x+\alpha_{i}$ is in $\hat{C}$ if and only if $(x+y)\to(x+y)+\alpha_{i}$ is in $\hat{C}$ for all $y\in L_{1}$.
Remark 1.15.
The purpose of periodicity is that the periodic cuts are precisely the ones that descend to $Q$. More precisely, let $\hat{C}$ be a cut on $\hat{Q}$, then for every arrow $a\in\hat{C}$, its $L_{1}$-orbit is contained in $\hat{C}$, and this orbit is naturally identified with a single arrow in $Q$. This clearly gives a bijection between $L_{1}$-periodic cuts of $\hat{Q}$ and cuts of $Q$, and to alleviate notation we identify these without explicit mention when no confusion is possible. Hence we also refer to the infinite cut quiver $\hat{Q}_{C}$ in analogy with 1.9.
Let us now characterise a higher preprojective cut purely in quiver terms. This has been proven explicitly for $n=2$ in [DramburgGasanova]. See also [HIO, Theorem 5.6] and [Giovannini, Corollary 4.7] for parts of this statement. The part of the proof showing finite-dimensionality uses a statement that is most conveniently proven using height functions, which we introduce in Section 2. The implication using height functions is in fact slightly stronger. We obtain that the cut $Q_{C}$ quiver is acyclic. In [HIO, Question 5.9], the authors ask whether the quiver of an $n$-hereditary algebra is acyclic, so our result confirms this for all algebras of type $\tilde{A}$, in any global dimension.
Proposition 1.16.
Let $C\subseteq Q_{1}$ be a cut. Then declaring $\deg(a)=1$ for all arrows $a\in C$ defines a higher preprojective structure on $kQ/I$ if and only if $C$ contains arrows of all types. In this case, the cut quiver $Q_{C}$ is acyclic.
Proof.
We need to check the conditions from 1.4. To do this, we use the bimodule resolution of $R\ast G\simeq kQ/I$ given by the superpotential $\omega$ as in [BSW]. In detail, $\omega$ is a linear combination of all elementary cycles in $Q$, and its $i$-th partial derivatives generate the $(n+1-i)$-th term in a projective bimodule resolution of $kQ/I$. Since we assumed that $C$ is a cut, it follows that with the grading induced from $C$, the superpotential $\omega$ becomes homogeneous of degree $1$. Now consider the last term in the projective bimodule resolution, which is generated by the 00-th derivatives of $\omega$. That means it is generated by $\omega$ itself, and hence in degree $1$ with respect to the induced grading. Therefore, every cut induces a grading of Gorenstein parameter $1$ on $kQ/I$. It remains to consider when the grading is locally finite dimensional. Since $kQ/I$ is generated by $Q$, and $C\subseteq Q_{1}$ is finite, it clearly suffices to check that the induced degree 00 part $(kQ/I)_{0}$ is finite dimensional. If $C$ does not contain arrows of all types, we assume without loss of generality that $C$ contains no arrows of type $1$. Then it follows from 1.12 that the sum of these arrows corresponds under the isomorphism $kQ/I\simeq R\ast G$ to the element $x_{1}\otimes 1$. Clearly, this element generates an infinite dimensional subalgebra isomorphic to $k[x_{1}]$ in $(kQ/I)_{0}$, so the cut is not higher preprojective. Now suppose that $C$ does contain arrows of all types. Then it follows from 2.14 that the quiver $Q_{C}$, which generates $(kQ/I)_{0}$, is acyclic and hence $(kQ/I)_{0}$ is finite dimensional. ∎
Note that the above proposition clearly applies to $L_{1}$-periodic cuts as well. We therefore now fix a cofinite sublattice $L_{1}\leq L_{0}$ and want to study $L_{1}$-periodic cuts of $\hat{Q}$. Once $L_{1}$ is fixed, we also fix the size of the quotient $|L_{0}/L_{1}|=|\hat{G}|=m$. Next, we recall the short exact sequence
$0\to L_{1}\xrightarrow[]{B^{\prime}}L_{0}\to\hat{G}\to 0$
Here, we choose $B^{\prime}$ as a matrix for the embedding, with respect to the fixed basis $\{\alpha_{1},\ldots,\alpha_{n}\}$ for $L_{0}$. Since $B$ is then defined up to $\operatorname{GL}_{n}(\mathbb{Z})$-right action, that means up to base change of $L_{1}$, we furthermore assume that $\det(B^{\prime})=|G|$, and consider $B^{\prime}$ up to $\operatorname{SL}_{n}(\mathbb{Z})$-right action. We will also refer to the matrix
$B=\left(\begin{array}[]{ c | c }B^{\prime}&\begin{matrix}1\\ \vdots\\ \end{matrix}\\ \hline\cr 0\cdots&1\end{array}\right).$
Remark 1.17.
It will be convenient at times to consider a specific $B^{\prime}$ which has only non-negative entries. This is possible by computing the Hermite normal form of $B^{\prime}$.
A crucial tool to study cuts are their types, which we define only for $Q$ but can be defined analogously for $\hat{Q}$ by choosing a fundamental domain for the $L_{1}$-action.
Definition 1.18.
For a cut $C$ on $Q$, we call
$\theta(C)=(\#\{a\in C\mid\theta(a)=i\})_{1\leq i\leq n+1}$
the type of $C$.
We make some remarks on these types.
Remark 1.19.
Let $C\subseteq Q_{1}$ be a cut of type $\theta(C)=(\gamma_{1},\ldots,\gamma_{n+1})$.
  1. (1)
    An easy counting argument shows that $\sum_{i=1}^{n+1}\gamma_{i}=|L_{0}/L_{1}|=|G|=m$.
  2. (2)
    From this, it follows that $0\leq\gamma_{i}\leq m$ for all $i$.
  3. (3)
    For every $i\in\{1,\ldots,n+1\}$, there is a trivial cut obtained by cutting all arrows of type $i$, which has type $(0,\ldots,0,m,0,\ldots,0)$.
It follows that the vectors $\theta(C)$ are integer points on the hyperplane $\sum_{i=1}^{n+1}x_{i}=m$, intersected with the cone spanned by trivial type vectors. This connection will be made more precise in Section 3.2 and Section 4.
2. Height functions
In this section, we introduce the main tool to investigate cuts, called the height function. This is inspired by the case $n=2$, where one can associate a height function to a perfect matching on a dimer model [IUDimerSpecial, kenyon2009lecturesdimers]. The same generalisation also appeared in [LammersDimer]. Since there is no dual bipartite graph in the case $n>2$, we instead define the height function directly on the lattice $L_{0}$. When $n=2$, this recovers a function which is equivalent to the height function for the hexagonal dimer model. As in the setup, we fix a cofinite $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$. Furthermore, we write $|L_{0}/L_{1}|=m$, and fix an $L_{1}$-periodic cut $C$.
Definition 2.1.
A map $h\colon L_{0}\to\mathbb{Z}$ is called a height function if it satisfies the following two conditions:
  1. (1)
    We have $h(0)=0$.
  2. (2)
    For any $x\in L_{0}$ and $\alpha_{i}\in\{\alpha_{1},\ldots,\alpha_{n+1}\}$ we have $h(x+\alpha_{i})\in\{h(x)+1,h(x)-n\}$.
If in addition $h(x+y)=h(x)+h(y)$ for all $x\in L_{0}$ and $y\in L_{1}$, we call $h$ an $L_{1}$-equivariant height function.
Next, we want to construct a height function for a (not necessarily $L_{1}$-periodic) cut $C$. To do this, we first consider the height increment of a path.
Definition 2.2.
Let $a\in\hat{Q}_{1}$ be an arrow. We define
$h_{C}(a)=\begin{cases}1&a\not\in C,\\ -n&a\in C.\end{cases}$
Let $p=a_{1}a_{2}\cdots a_{l}$ be a path in $\hat{Q}$. Then we define the height increment along $p$ as
$h_{C}(p)=\sum_{i=1}^{l}h_{C}(a_{i}).$
Remark 2.3.
For any pair of points $x,y\in L_{0}$ there is a path from $x$ to $y$ in $\hat{Q}$. In other words, $\hat{Q}$ is strongly connected. This follows from the more general observation that $Q$ as an undirected graph $\hat{Q}$ is connected, and that any arrow can be completed into a cycle. Choose an undirected path from $x$ to $y$. If this undirected path contains a reversed arrow, we can simply replace this arrow by a directed path that completes the arrow into a cycle.
In the next proposition, we show that the height increment of any closed path is 00. We do this by showing that we may deform a path by commuting arrows of different types or contracting elementary cycles without affecting its height increment. As a consequence, we will obtain that the height increments of any two parallel paths is the same.
Proposition 2.4.
Let $c\colon x\to\ldots\to x$ be a cycle in $\hat{Q}$. Then $h_{C}(c)=0$.
Proof.
First, note that any elementary cycle $c$ has length $n+1$, and exactly one of its arrows is in $C$, so we have $h_{C}(c)=0$. Next, let $p_{1}=a_{i}a_{j}$ and $p_{2}=a_{j}a_{i}$ be two paths from $x$ to $x+\alpha_{i}+\alpha_{j}$. Then $h_{C}(p_{1})=h_{C}(p_{2})$. Indeed, this is obvious for $i=j$, and otherwise $p_{1}$ and $p_{2}$ can be completed into elementary cycles by the same path $p_{3}$, that is, there exists a path $p_{3}$ such that $p_{1}p_{3}$ and $p_{2}p_{3}$ are elementary cycles. But that means
$h_{C}(p_{1})+h_{C}(p_{3})=h_{C}(p_{1}p_{3})=0=h_{C}(p_{2}p_{3})=h_{C}(p_{2})+h_{C}(p_{3}),$
and hence $h_{C}(p_{1})=h_{C}(p_{2})$. This means that deforming a path using commutativity relations $a_{i}a_{j}=a_{j}a_{i}$ does not change the height increment. Finally, let $c$ be any cycle $x\to\ldots\to x$ and let $t_{i}$ denote the number of arrows of type $i$ in $c$ for all $i\in\{1,\ldots,n+1\}$. Then
$0=\sum_{i=1}^{n+1}t_{i}\alpha_{i}=\sum_{i=1}^{n}t_{i}\alpha_{i}+t_{n+1}\left(-\sum_{i=1}^{n}\alpha_{i}\right)=\sum_{i=1}^{n}(t_{i}-t_{n+1})\alpha_{i}.$
from which we conclude that $t_{1}=t_{2}=\ldots=t_{n+1}$. Using commutativity relations, this path can therefore be deformed into the path $c^{t_{1}}$, where $c$ is an elementary cycle. Thus, $c$ and its powers give a height increment 00. ∎
Corollary 2.5.
Let $p$, $q$ be paths from $x$ to $y$ in $\hat{Q}$. Then
$h_{C}(p)=h_{C}(q).$
Proof.
Let $r\colon y\to\ldots\to x$ be any path in $\hat{Q}$, which we know exists by 2.3. Then both $pr$ and $qr$ are cycles $x\to\ldots\to x$. Then
$h_{C}(p)+h_{C}(r)=h_{C}(pr)=0=h_{C}(qr)=h_{C}(q)+h_{C}(r),$
which implies $h_{C}(p)=h_{C}(q)$. ∎
This allows us to make the following construction.
Proposition 2.6.
Let $C$ be a cut, and for $x\in L_{0}$ denote by $p_{x}$ any path from 00 to $x$ in $\hat{Q}$. Then the map
$h_{C}\colon L_{0}\to\mathbb{Z},\ h_{C}(x)=h_{C}(p_{x})$
is a height function. Furthermore, the map $C\mapsto h_{C}$ is a bijection between the set of cuts and the set of height functions, which restricts to a bijection between $L_{1}$-periodic cuts and $L_{1}$-equivariant height functions. Its inverse is given by
$h\mapsto C_{h}=\{(x\to x+\alpha_{i})\in\hat{Q}_{1}\mid h(x+\alpha_{i})=h(x)-n\}.$
Proof.
It follows from 2.5 that $h_{C}$ is well defined, since any two paths $p$, $q$ from 00 to $x$ have the same height increment. Next, recall that for every arrow $a\in\hat{Q}$, we have $h_{C}(a)\in\{1,-n\}$, and hence $h_{C}(x+\alpha_{i})\in\{h_{C}(x)+1,h_{C}(x)-n\}$ for any $x\in L_{0}$ and any $i\in\{1,\ldots,n+1\}$. Since any path from 00 to 00 is a cycle, it follows that $h_{C}(0)=0$, so $h_{C}$ is indeed a height function. To prove the bijection, consider for any height function $h$ the set
$C_{h}=\{(x\to x+\alpha_{i})\in\hat{Q}_{1}\mid h(x+\alpha_{i})=h(x)-n\}.$
It follows that for every elementary cycle there is exactly one arrow in the cycle so that $h$ decreases along this arrow, and hence $C_{h}$ is a cut. Thus, $C\mapsto h_{C}$ and $h\mapsto C_{h}$ are mututally inverse bijections between the set of all cuts and all height functions, which clearly restrict to bijections between periodic cuts and equivariant height functions. ∎
Just as with cuts, we will often drop the adjective $L_{1}$-equivariant for a height function, since we will only consider those which are $L_{1}$-equivariant.
Remark 2.7.
Note that a height function $h$ does not descend to a function on $L_{0}/L_{1}$, but equivariance implies that a height function $h$ restricts to a homomorphism $h\colon L_{1}\to\mathbb{Z}$.
Next, we aim to compute the values of the homomorphism $h_{C}$ on $L_{1}$ explicitly, based on the type of $C$. For the computation, we fix the type $\theta(C)=\gamma=(\gamma_{i})_{1\leq i\leq n+1}$. Due to the discrepancy between the number of entries in $\gamma$ and the dimension of $L_{0}$, the following notation will be convenient throughout.
Definition 2.8.
  1. (1)
[…]
  2. (2)
    For row or column vectors $x=(x_{1},\ldots,x_{l_{1}})$ and $y=(y_{1},\ldots,y_{l_{2}})$, where $l_{1},l_{2}\in\{n,n+1\}$, we write
    $\langle x,y\rangle_{n}=\sum_{i=1}^{n}x_{i}y_{i}.$
    If $l_{1}=l_{2}=n+1$, we also write
    $\langle x,y\rangle_{n+1}=\sum_{i=1}^{n+1}x_{i}y_{i}.$
Proposition 2.9.
For any $y\in L_{1}$ and $L_{1}$-periodic cut $C$ of type $\gamma=(\gamma_{1},\ldots,\gamma_{n+1})$ we have
$h_{C}(y)=\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}.$
Proof.
Let $o_{i}$ denote the order of $\alpha_{i}$ in $L_{0}/L_{1}$. First we prove the claim for the points $o_{i}\alpha_{i}\in L_{1}$. To compute $h_{C}(o_{i}\alpha_{i})$, consider the path $0\to\alpha_{i}\to 2\alpha_{i}\to\cdots\to o_{i}\alpha_{i}$, where each arrow is of type $i$. Then $h_{C}(o_{i}\alpha_{i})=o_{i}-(n+1)\theta_{i}^{\prime}$, where $\theta_{i}^{\prime}=\#\{j\in[0,o_{i})\mid(j\alpha_{i}\to(j+1)\alpha_{i})\in C\}$ is the number of arrows in this path which lie in $C$. Next, consider any path $x\to x+\alpha_{i}\to\ldots\to x+o_{i}\alpha_{i}$. Since $o_{i}\alpha_{i}\in L_{1}$, equivariance implies
$h_{C}(x+o_{i}\alpha_{i})-h_{C}(x)=h_{C}(o_{i}\alpha_{i})=o_{i}-(n+1)\theta_{i}^{\prime}.$
Equivalently, this means that the number of arrows in $C$ in this path is also
$\theta_{i}^{\prime}=\frac{1}{n+1}(o_{i}-h_{C}(o_{i}\alpha_{i})).$
To cover all arrows of type $i$ in $Q$ exactly once by these paths, we let $\{x_{1},\ldots,x_{\frac{m}{o_{i}}}\}$ be a system of representatives of the cosets of $\langle\alpha_{i}\rangle$ in $L_{0}/L_{1}$. This means that each arrow of type $i$ in $Q$ lies in exactly one cycle
$x_{j}+L_{1}\to x_{j}+\alpha_{i}+L_{1}\to\cdots\to x_{j}+o_{i}\alpha_{i}+L_{1},$
and hence we have that
$\gamma_{i}=\frac{m}{o_{i}}\cdot\theta_{i}^{\prime}=\frac{m}{o_{i}}\frac{1}{n+1}(o_{i}-h_{C}(o_{i}\alpha_{i})),$
which gives
$h_{C}(o_{i}\alpha_{i})=o_{i}-\frac{n+1}{m}\gamma_{i}o_{i}=o_{i}\left(1-\frac{n+1}{m}\gamma_{i}\right),$
which is the desired statement, given that $o_{i}\alpha_{i}=(0,\ldots,0,o_{i},0,\ldots,0)^{\top}$. Now for an arbitrary $y=(y_{1},\ldots,y_{n})^{\top}\in L_{1}$ we compute $h_{C}(my)$ in two different ways. On the one hand we have
$h_{C}(my)=h_{C}\left(\sum_{i=1}^{n}{my_{i}\alpha_{i}}\right)=h_{C}\left(\sum_{i=1}^{n}{y_{i}\cdot\frac{m}{o_{i}}\cdot(o_{i}\alpha_{i})}\right)=\sum_{i=1}^{n}{y_{i}\cdot\frac{m}{o_{i}}\cdot h_{C}(o_{i}\alpha_{i})},$
since $o_{i}\alpha_{i}\in L_{1}$ and since $\frac{m}{o_{i}}\in\mathbb{Z}$ for all $i$. Since we already established the height values of $o_{i}a_{i}$, the above equality becomes
$h_{C}(my)$ $=\sum_{i=1}^{n}{y_{i}\cdot\frac{m}{o_{i}}\cdot o_{i}\left(1-\frac{n+1}{m}\gamma_{i}\right)}$
$=m\sum_{i=1}^{n}{y_{i}\cdot\left(1-\frac{n+1}{m}\gamma_{i}\right)}=m\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}$
On the other hand, $h_{C}(my)=mh_{C}(y)$ since $y\in L_{1}$. This completes the proof. ∎
The homomorphism from the above proposition will be useful later on, so we fix notation for it here.
Definition 2.10.
For a type $\gamma$ we write
$h_{\gamma}\colon L_{1}\to\mathbb{Z},y\mapsto h_{C}(y)$
where $C$ is an arbitrary cut of type $\gamma$.
The formula we give above for $h_{\gamma}$ is invertible, so we can use $h_{\gamma}$ or the type $\gamma$ interchangeably, as the following proposition shows.
Proposition 2.11.
Let $\gamma$ and $\delta$ be types of $L_{1}$-periodic cuts. Then $h_{\gamma}=h_{\delta}$ if and only if $\gamma=\delta$.
Proof.
One implication is obvious. For the other, assume that $h_{\delta}(y)=h_{\gamma}(y)$ for all $y\in L_{1}$. Therefore, from 2.9 we get
$\langle y,\textbf{1}\rangle_{n}-(n+1)\frac{\langle y,\gamma\rangle_{n}}{m}=h_{\gamma}(y)=h_{\delta}(y)=\langle y,\textbf{1}\rangle_{n}-(n+1)\frac{\langle y,\delta\rangle_{n}}{m},$
which implies $\langle y,\gamma-\delta\rangle_{n}=0$ for all $y\in L_{1}$. Setting $y=m\alpha_{i}\in L_{1}$ for all $i\in\{1,\ldots,n\}$, we obtain $\gamma_{i}=\delta_{i}$ for all $i\in\{1,\ldots,n\}$. Since $\langle\gamma,\textbf{1}\rangle_{n+1}=m=\langle\delta,\textbf{1}\rangle_{n+1}$, we also get $\gamma_{n+1}=\delta_{n+1}$. ∎
The following reformulation of 2.6,2.9 and 2.11 will be useful in Section 5
Corollary 2.12.
Let $h$ be an $L_{1}$-equivariant height function. Then $h(y)=\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}$ for all $y\in L_{1}$ if and only if $C_{h}$ has type $\gamma$.
We are ready to show the necessity of the condition from 3.5.
Proposition 2.13.
Let $C$ be an $L_{1}$-periodic cut on $\hat{Q}$ of type $\gamma$. Then we have
$\gamma\cdot B\in m\mathbb{Z}^{1\times(n+1)}$
Proof.
Write $\gamma=(\gamma_{1},\ldots,\gamma_{n+1})$. It suffices to show that
$(\gamma_{1},\ldots,\gamma_{n})y\in m\mathbb{Z}^{1\times n}$
for all columns $y$ of $B^{\prime}$. Recall from 1.17 that we may choose $B^{\prime}$ with non-negative entries. It suffices to show the claim for this choice of $B^{\prime}$ since clearly the $\operatorname{SL}_{n}(\mathbb{Z})$-action leaves invariant the sublattice $m\mathbb{Z}^{1\times n}$, more precisely we have $m\mathbb{Z}^{1\times n}\cdot\operatorname{SL}_{n}(\mathbb{Z})=m\mathbb{Z}^{1\times n}$. Thus, let $y=(y_{1},\ldots,y_{n})^{\top}$ be a column, where $y_{i}\geq 0$. We follow the path
$0\to\alpha_{1}\to 2\alpha_{1}\to\cdots\to y_{1}\alpha_{1}\to y_{1}\alpha_{1}+\alpha_{2}\to\cdots\to y_{1}\alpha_{1}+y_{2}\alpha_{2}\to\cdots\to y.$
This path has $\langle y,\textbf{1}\rangle_{n}$ arrows. Denote by $\varepsilon$ the number of arrows in this path that belong to $C$. Then, $h_{C}(y)=\langle y,\textbf{1}\rangle_{n}-(n+1)\varepsilon$. However, it follows from 2.9 that
$h_{C}(y)=\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}=\langle y,\textbf{1}\rangle_{n}-\frac{n+1}{m}\langle y,\gamma\rangle_{n}.$
Comparing the two expressions for $h_{C}(y)$, we obtain $\langle y,\gamma\rangle_{n}=m\varepsilon\in m\mathbb{Z}$, which proves our claim. ∎
We conclude this section by using height functions to prove that cuts of positive type give acyclic cut quivers. In other words, a cut of positive type is higher preprojective, see 1.16.
Lemma 2.14.
Let $C$ be a cut of type $\theta(C)=\gamma=(\gamma_{1},\gamma_{2},\ldots,\gamma_{n+1})$ such that $\gamma_{i}>0$ for all $i$. Then there is no infinite path in $\hat{Q}_{C}$. Equivalently, the cut quiver $Q_{C}$ for $L_{0}/L_{1}$ is acyclic.
Proof.
Assume there is an infinite path in $\hat{Q}_{C}$. Since $L_{0}/L_{1}$ is a finite group, there exists a non-empty subpath $c\colon x\to\ldots\to x+y$ in $\hat{Q}_{C}$ such that $y\in L_{1}$. Let $y=(y_{1},\ldots,y_{n})^{\top}$. We know that $h_{C}(x+y)-h_{C}(x)=h_{C}(y)$, therefore, the height increment of $c$ is $h_{C}(c)=h_{C}(y)=h_{\gamma}(y)$. Now, the height increases exactly by $1$ each step and hence $c$ has exactly $h_{\gamma}(y)$ steps. Let $t=(t_{1},\ldots,t_{n+1})$, where $t_{i}$ denotes the number of arrows of type $i$ in $c$. Then we have $t_{n+1}=h_{\gamma}(y)-\langle t,\textbf{1}\rangle_{n}$ and
$\sum_{i=1}^{n}{y_{i}\alpha_{i}}$ $=\sum_{i=1}^{n+1}{t_{i}\alpha_{i}}$
$=\sum_{i=1}^{n}{t_{i}\alpha_{i}}+(h_{\gamma}(y)-\langle t,\textbf{1}\rangle_{n})\alpha_{n+1}$
$=\sum_{i=1}^{n}{t_{i}\alpha_{i}}+(h_{\gamma}(y)-\langle t,\textbf{1}\rangle_{n})\left(-\sum_{i=1}^{n}{\alpha_{i}}\right)$
$=\sum_{i=1}^{n}{t_{i}\alpha_{i}}+(\langle t,\textbf{1}\rangle_{n}-h_{\gamma}(y))\left(\sum_{i=1}^{n}{\alpha_{i}}\right)$
$=\sum_{i=1}^{n}{\alpha_{i}(\langle t,\textbf{1}\rangle_{n}+t_{i}-h_{\gamma}(y))}.$
Comparing the $\alpha_{i}$-coefficients, we obtain for each $i\in\{1,\ldots,n\}$ that
$y_{i}=\langle t,\textbf{1}\rangle_{n}+t_{i}-h_{\gamma}(y).$
Summing over all $i$, we obtain
$\langle y,\textbf{1}\rangle_{n}=n\langle t,\textbf{1}\rangle_{n}+\langle t,\textbf{1}\rangle_{n}-nh_{\gamma}(y)=(n+1)\langle t,\textbf{1}\rangle_{n}-nh_{\gamma}(y),$
in other words,
$\langle t,\textbf{1}\rangle_{n}=\frac{\langle y,\textbf{1}\rangle_{n}+nh_{\gamma}(y)}{n+1}.$
This, combined with the equation for the $\alpha_{i}$-coefficient, gives
$t_{i}=y_{i}+h_{\gamma}(y)-\langle t,\textbf{1}\rangle_{n}=y_{i}+h_{\gamma}(y)-\frac{\langle y,\textbf{1}\rangle_{n}+nh_{\gamma}(y)}{n+1}=y_{i}+\frac{h_{\gamma}(y)-\langle y,\textbf{1}\rangle_{n}}{n+1}$
for all $i\in\{1,\ldots,n\}$. We also conclude that
$t_{n+1}=h_{\gamma}(y)-\langle t,\textbf{1}\rangle_{n}=h_{\gamma}(y)-\frac{\langle y,\textbf{1}\rangle_{n}+nh_{\gamma}(y)}{n+1}=\frac{h_{\gamma}(y)-\langle y,\textbf{1}\rangle_{n}}{n+1}.$
Now,
$\langle t,\gamma\rangle_{n+1}$ $=\langle y,\gamma\rangle_{n}+\frac{h_{\gamma}(y)-\langle y,\textbf{1}\rangle_{n}}{n+1}\langle\gamma,\textbf{1}\rangle_{n+1}$
$=\langle y,\gamma\rangle_{n}+\frac{h_{\gamma}(y)-\langle y,\textbf{1}\rangle_{n}}{n+1}\cdot m$
$=\frac{m}{n+1}\left(\left\langle y,\frac{n+1}{m}\gamma\right\rangle_{n}+h_{\gamma}(y)-\langle y,\textbf{1}\rangle_{n}\right)$
$=\frac{m}{n+1}\left(h_{\gamma}(y)-\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}\right)=0,$
where the last equality follows from 2.9. Since $\gamma_{1},\ldots,\gamma_{n+1}>0$ and $t_{1},\ldots,t_{n+1}\geq 0$, we conclude $t_{1}=\ldots=t_{n+1}=0$, which implies that the path $c$ is empty, which contradicts our assumption. ∎
3. Construction of cuts and simplices of types
3.1. Construction of cuts
In this subsection, we construct a cut for a given type, following the ideas outlined in [DramburgGasanova]. We fix as before $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$ and $m=|L_{0}/L_{1}|$. We begin by recalling the construction by Amiot-Iyama-Reiten from [AIR, Section 5], which produces a cut for the quiver $Q$ of a cyclic group generated by a junior element. In the following, recall that we refer by $x\bmod m$ to the smallest non-negative representative of the equivalence class $x+m\mathbb{Z}$.
Proposition 3.1.
[AIR, Theorem 5.6] Suppose $L_{0}/L_{1}$ is cyclic. If there exist $e_{1},\ldots,e_{n+1}\in\{0,\ldots,m-1\}$ with $\sum_{i=1}^{n+1}e_{i}=m$ and an isomorphism
$\xi\colon L_{0}/L_{1}\to\mathbb{Z}/m\mathbb{Z}$
such that $\xi(\alpha_{i}+L_{1})=e_{i}+m\mathbb{Z}$ for all $i\in\{1,\ldots,n+1\}$, then there exists a cut of type $(e_{1},\ldots,e_{n+1})$ on $Q$.
Proof.
We use the isomorphism $\xi$ to label the vertices of $Q$ by $\{0,\ldots,m-1\}$, by identifying $\xi(x)$ with its smallest non-negative representative. This way, every arrow $x\to x+\alpha_{i}$ of type $i$ corresponds to $j\to(j+e_{i})\bmod m$ for $\xi(x)=j\in\{0,\ldots,m-1\}$. We then define
$C=\{j\to(j+e_{i})\bmod m\mid j>(j+e_{i})\bmod m\}.$
We claim that $C$ is a cut. To see this, consider an elementary cycle $c$. Let $D$ denote the set of decreasing arrows of $c$, that is, $D=C\cap c$, and let $N=c\setminus D$ be the set of non-decreasing arrows in $c$. Each decreasing arrow acts by adding $e_{i}-m$, and any non-decreasing arrow acts by adding $e_{i}$. Then we get
00 0 $=\sum_{a\in N}e_{\theta(a)}+\sum_{a\in D}{(e_{\theta(a)}-m)}$
$=\sum_{a\in c}{e_{\theta(a)}}-\sum_{a\in D}{m}=m-\sum_{a\in D}{m},$
which means there is exactly one decreasing arrow in $c$. To see that $C$ is of the correct type, simply note that an arrow of type $i$ is in $C$ if and only if it ends at one of the vertices $10,1,\ldots,e_{i}-$ 1. ∎
Remark 3.2.
Note that the condition on the $e_{i}$ in 3.1 means that we are considering the group $\mathbb{Z}/m\mathbb{Z}$ as generated inside of $\operatorname{SL}_{n+1}(k)$ by the element $\frac{1}{m}(e_{1},\ldots,e_{n+1})$. In the terminology of Ito-Reid [ItoReid], this is a junior element.
Now, our construction proceeds in two steps. We construct an epimorphism $\xi$ as above for each potential type $\gamma$, and then lift a cut from this to our quiver. Equivalently, we are showing that we can find a cyclic subgroup of $G$ generated by the correct junior to realise $\gamma$ as a type, and extend to all of $G$.
Lemma 3.3.
Let $\gamma=(\gamma_{1},\ldots,\gamma_{n+1})\in\mathbb{N}^{1\times(n+1)}$ such that $\langle\gamma,\mathbf{1}\rangle_{n+1}=m$ and $\gamma B\in m\mathbb{Z}^{1\times(n+1)}$. Then
$\xi_{\gamma}\colon L_{0}\to\mathbb{Z}/m\mathbb{Z},x\to\langle x,\gamma\rangle_{n}+m\mathbb{Z}$
is a group homomorphism with $\operatorname{Ker}(\xi_{\gamma})\supseteq L_{1}$ and cyclic image of order $\frac{m}{\gcd(\gamma_{1},\ldots,\gamma_{n+1})}$.
Proof.
It is clear that $\xi_{\gamma}$ is a group homomorphism. We first investigate
$\operatorname{Ker}(\xi_{\gamma})=\{x\in L_{0}\mid\langle x,\gamma\rangle_{n}\equiv_{m}0\}.$
Recall that any $y\in L_{1}$ can be written as $y=B^{\prime}v$ for some $v\in\mathbb{Z}^{n}$. Using our assumption that $\gamma B\in m\mathbb{Z}^{1\times(n+1)}$, we find
$\xi_{\gamma}(y)=\langle\gamma,B^{\prime}v\rangle_{n}=\langle\gamma B,v\rangle_{n}\equiv_{m}0,$
and therefore $\operatorname{Ker}(\xi_{\gamma})\supseteq L_{1}$. To compute the order of $\operatorname{Im}(\xi_{\gamma})$, note that we can write $d=\gcd(\gamma_{1},\ldots,\gamma_{n+1})=\gcd(\gamma_{1},\ldots,\gamma_{n},m)$ as an integer combination of $\gamma_{1}\ldots,\gamma_{n}$ and $m$. That means there exist numbers $x_{1},\ldots,x_{n}\in\mathbb{Z}$ such that $\xi_{\gamma}(x_{1},\ldots,x_{n})=d+m\mathbb{Z}$. Since $d\mid\gamma_{1},\ldots,\gamma_{n}$, it follows that the subgroup $\operatorname{Im}(\xi_{\gamma})\subseteq\mathbb{Z}/m\mathbb{Z}$ is generated by the coset of $d$, and therefore has order $\frac{m}{\gcd(\gamma_{1},\ldots,\gamma_{n+1})}$. ∎
Proposition 3.4.
Let $\gamma$ be as in 3.3 and let $L_{2}=\operatorname{Ker}(\xi_{\gamma})$. Then there exists an $L_{2}$-periodic cut $C$ of type $(\frac{\gamma_{1}}{d},\ldots,\frac{\gamma_{n+1}}{d})$, where $d=\gcd(\gamma_{1},\ldots,\gamma_{n+1})$. In particular, $C$ is $L_{1}$-periodic of type $(\gamma_{1},\ldots\gamma_{n+1})$.
Proof.
We write $m^{\prime}=\frac{m}{d}$. If $m^{\prime}=1$, one of the $\gamma_{i}$ is $m$ and $L_{2}=L_{0}$, hence we take $C$ to be the constant cut of the corresponding type. Otherwise, by 3.3, the morphism $\xi_{\gamma}$ has a cyclic image $\operatorname{Im}(\xi_{\gamma})\simeq\mathbb{Z}/m^{\prime}\mathbb{Z}$. Hence, we have that $L_{0}/L_{2}$ is cyclic of order $m^{\prime}$, and we will use 3.1 to construct an $L_{2}$-periodic cut. To do this, we factor $\xi_{\gamma}$ as the projection $q\colon L_{0}\to L_{0}/L_{2}$, followed by the induced isomorphism $\overline{\xi_{\gamma}}$ to $\mathbb{Z}/m^{\prime}\mathbb{Z}$, and the inclusion of $\mathbb{Z}/m^{\prime}\mathbb{Z}$ as $\operatorname{Im}(\xi_{\gamma})\leq\mathbb{Z}/m\mathbb{Z}$.
[…]
We claim that $\overline{\xi_{\gamma}}\circ q$ satisfies the conditions from 3.1. Since $\xi_{\gamma}(\alpha_{i})=\gamma_{i}+m\mathbb{Z}$, we have that $(\overline{\xi_{\gamma}}\circ q)(\alpha_{i})=\frac{\gamma_{i}}{d}+m\mathbb{Z}$. Furthermore, since $\langle\gamma,\mathbf{1}\rangle_{n+1}=m$, it follows that $\langle\frac{1}{d}\gamma,\mathbf{1}\rangle_{n+1}=\frac{m}{d}=m^{\prime}$. Therefore, the values $e_{i}=\frac{\gamma_{i}}{d}$ together with $\overline{\xi_{\gamma}}\circ q$ satisfy the conditions from 3.1. Thus, there exists an $L_{2}$-periodic cut $C$ of type $\frac{1}{d}\gamma$, and since $L_{0}/L_{2}$ has index $d$ in $L_{0}/L_{1}$, it follows that the type of $C$ as an $L_{1}$-periodic cut is $\gamma$. ∎
This leads to our first main theorem.
Theorem 3.5.
Let $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$ be a cofinite embedding of lattices and $m=|L_{0}/L_{1}|$. Then a vector $\gamma\in\mathbb{N}^{1\times(n+1)}$ is a type of a cut if and only if $\langle\gamma,\mathbf{1}\rangle_{n+1}=m$ and
$\gamma B\in m\mathbb{Z}^{1\times(n+1)}.$
Furthermore, $\gamma$ is a type of a higher preprojective cut if and only if in addition, $\gamma_{i}>0$ for all $i\in\{1,\ldots,n+1\}$.
Proof.
The first implication follows from 2.13. The other implication follows from 3.4. The statement about higher preprojectivity is 1.16. ∎
3.2. Simplices of types
Now, we want to investigate the problem of when a higher preprojective cut exists for a given $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$. To do this, it suffices to show that a type of a higher preprojective cut exists. We first note, as alluded to in 1.19, that the set of all types of cuts forms a lattice simplex.
Remark 3.6.
Let $\gamma$ be the type of a cut. Then we can rewrite 3.5 as saying that $\gamma$ lies in a lattice $N=m\mathbb{Z}^{1\times(n+1)}B^{-1}$. Combining this with the condition that $\gamma_{i}\geq 0$ for all $i\in\{1,\ldots,n+1\}$ and $\langle\gamma,\mathbf{1}\rangle_{n+1}=m$, we see that $\gamma$ lies in the convex hull $P=\operatorname{Conv}(\{\theta(C)\mid C\text{ is a trivial cut}\})$. Recall that a trivial cut $C$ is given by placing all arrows of a fixed type $i$ in degree $1$, and that its type therefore is given by $\theta(C)_{i}=m$. We fix this polytope $P$, and since the vertices of $P$ are in $N$ we consider it as a lattice polytope. Furthermore, since $P$ has $n+1$ vertices, it is a lattice simplex. We also note that an isomorphic lattice polytope in the standard $\mathbb{Z}^{n+1}$ lattice is given by the convex hull $\operatorname{Conv}(\operatorname{Cols}(B^{\top}))$ of the rows of $B$.
Corollary 3.7.
The skew-group algebra $R\ast G$ admits a higher preprojective cut if and only if the polytope $P$ has an internal lattice point.
Definition 3.8.
A lattice polytope $P$ is called hollow if there exist no interior lattice points in $P$. A lattice polytope is called empty if the only lattice points in $P$ are the vertices.
Given a lattice polytope, it is of course possible to check whether it contains an interior lattice point. However, given just the group $G$, we are not aware of an easy criterion for the existence of a higher preprojective cut that avoids computing $P$. Let us comment on this problem for cyclic groups.
Remark 3.9.
Consider a cyclic group $G=\langle\frac{1}{m}(e_{1},\ldots,e_{n+1})\rangle\leq\operatorname{SL}_{n+1}(k)$, and fix some $B$ for it. Consider any $\gamma\in\mathbb{Z}^{1\times(n+1)}$ such that
$\gamma B\in m\mathbb{Z}^{1\times(n+1)}$
and $0\leq\gamma_{i}<m$ for all $i$. Then $\frac{1}{m}(\gamma_{1},\ldots,\gamma_{n+1})$ defines an element in $G$, and conversely every element $\frac{1}{m}(f_{1},\ldots,f_{n+1})$ in $G$ satisfies
$(f_{1},\ldots,f_{n+1})B\in m\mathbb{Z}^{1\times(n+1)}.$
Hence, to find types of cuts, we need to find elements $\frac{1}{m}(f_{1},\ldots,f_{n+1})$ in $G$ such that $\sum_{i=1}^{n+1}f_{i}=m$. These are simply the junior elements (see Section 4), so we need to compute all elements in $G$ to decide whether $G$ has a junior element $\frac{1}{m}(f_{1},\ldots,f_{n+1})$ with all $f_{i}\neq 0$, and hence whether a higher preprojective cut exists. We are not aware of a general procedure that avoids computing all elements of $G$ to check whether such an element, respectively such an internal point in $P$, exists. In [DramburgGasanova, Example 6.5], an example for the group of order $11$ was given, further illustrating this point.
The upshot of this perspective is that a generic lattice polytope is not hollow, as the following theorem by Nill and Ziegler shows. To phrase it, we recall that a lattice projection is a surjective affine map
$f\colon\mathbb{R}^{n}\twoheadrightarrow\mathbb{R}^{m},x\mapsto Ax+b$
which preserves the lattices, that means which satisfies that $f(\mathbb{Z}^{n})=\mathbb{Z}^{m}$. In this case, the lattice projection of the lattice polytope $P\subseteq\mathbb{R}^{n}$ is $f(P)$, which is a lattice polytope in $\mathbb{R}^{m}$.
Theorem 3.10.
[NillZiegler, Theorem 1.1] For every $n\in\mathbb{N}$, there are only finitely many hollow $n$-dimensional lattice polytopes that do not project onto a lower dimensional hollow lattice polytope.
The finitely many exceptions in the above theorem are called exceptional polytopes.
Corollary 3.11.
Let $G\leq\operatorname{SL}_{n+1}(k)$ be finite abelian such that the skew-group algebra $R\ast G$ does not afford a higher preprojective cut. Let $P$ be the simplex associated to $R\ast G$.
  1. (1)
    If $P$ is an exceptional polytope, the order of $G$ is bounded.
  2. (2)
    For all but finitely many $G$, the polytope $P$ admits a lower dimensional hollow lattice projection.
Proof.
To prove the first statement, simply note that the proof of [NillZiegler, Theorem 1.1] shows that the volume of an exceptional hollow polytope is bounded and hence the order of an abelian group $G\leq\operatorname{SL}_{n+1}(k)$ giving rise to such an exceptional polytope is bounded as well. The second statement follows directly from 3.7 and 3.10. ∎
Remark 3.12.
The width (see e.g. [NillZiegler, Section 1.4]) of a lattice polytope does not decrease under lattice projection. It therefore follows from 3.10 that the width of a hollow lattice polytope of a fixed dimension is bounded. However, the width of the simplex $P$ of types is not bounded, so in this sense a generic simplex we consider has an internal lattice point, which means that generically the algebra $kQ/I$ we consider admits a higher preprojective cut.
The hollow lattice polytopes in dimension $2$ are known, and lead to the classification of $3$-preprojective algebras of type $\tilde{A}$ as given in [DramburgGasanova]. In higher dimensions, we are not aware of a complete classification of the exceptional hollow polytopes. Let us conclude by comparing the infinite families from 3.10 and from Thibault’s criterion 1.7.
Remark 3.13.
Let $G\leq\operatorname{SL}_{n+1}(k)$ be abelian such that the embedding of $G$ factors through some embedding of $\operatorname{SL}_{n_{1}}(k)\times\operatorname{SL}_{n_{2}}(k)$ with $n_{1}+n_{2}=n+1$ and $n_{1},n_{2}>0$. Then it follows from 1.7 that the simplex $P$ for $G$ is hollow, and one can show that $P$ projects onto a lower dimensional hollow lattice polytope. However, this is not an equivalence: The group $G\simeq C_{2}\times C_{2}\times C_{2}$ embeds into $\operatorname{SL}_{4}(k)$ in a unique way so that the embedding does factor through a product $\operatorname{SL}_{n_{1}}(k)\times\operatorname{SL}_{n_{2}}(k)$. The resulting simplex $P$ of types of cuts projects onto the simplex $P^{\prime}$ of types of cuts for the group $C_{2}\times C_{2}$ embedded in $\operatorname{SL}_{3}(k)$. It is easy to compute that $P^{\prime}$ is the unique exceptional hollow lattice triangle. It follows from 3.10 that $P$ is hollow, but the embedding of $G$ does not factor.
We therefore note that it would be interesting to find a more refined version of 1.7 that covers more of the infinite families of hollow polytopes. Ideally, this would give a purely algebraic description of the polytopes that appear in the second part of 3.11. In particular, the orders of the lattice generators $\alpha_{1},\ldots,\alpha_{n+1}$ modulo $L_{1}$ may be used to describe the width of the lattice simplex. Since the width of a hollow lattice polytope is bounded, this might be used to characterise these infinite families of hollow simplices.
4. Toric geometry
In this section, we want to highlight the connections between higher preprojective cuts and toric geometry. When $n=1$, these observations are part of the classical McKay correspondence. When $n=2$, dimer models can be used and these observations were made by Nakajima in [Nakajima], based on the work of Ishii and Ueda [IUDimerSpecial, IUDimerCrepant, IUModuliDimer]. We show that the same correspondence holds for any $n$: The lattice simplex $P$ of types of cuts is naturally isomorphic to the lattice simplex of junior elements in $G$, which in turn is the slice of the cone of the toric variety $X=\operatorname{Spec}(R^{G})$ at height $1$. For a general resource on the parts of toric geometry we use, we refer the reader to [CLS, Chapter 1], and we will use the standard notation established in this book. In particular, that means that when we consider an affine toric variety $X$, we denote by $N$ the lattice of $1$-parameter subgroups and by $C\subseteq N_{\mathbb{R}}$ the cone defining $X$. As before, we assume that $G\leq\operatorname{SL}_{n+1}(k)$ is abelian and acts on the polynomial ring $R=k[x_{1},\ldots,x_{n+1}]$. We assume without loss of generality that the action of $G$ is diagonalised, so that the invariant ring $R^{G}$ is generated by monomials. We begin with the following observation.
Remark 4.1.
The invariant ring $R^{G}$ embeds into $R\ast G$ via $r\to r\otimes 1$, and in this way we obtain precisely the center of the skew-group algebra $Z(R\ast G)=R^{G}$, which is true in more generality for any commutative domain $R$ and finite group $G$ [MarcosMartinezVMartins, Proposition 4.8]. Since $R^{G}$ is generated by monomials, it is the coordinate ring of an affine toric variety. The torus action on $X=\operatorname{Spec}(R^{G})$ is diagonal in the coordinates induced from the invariant monomials. We fix these coordinates.
Next, we note that the center is automatically a graded subalgebra.
Remark 4.2.
Let $\Gamma=\bigoplus_{i\in\mathbb{Z}}\Gamma_{i}$ be a $\mathbb{Z}$-graded algebra, and let $z\in Z(\Gamma)$ be a central element. This is equivalent to $az=za$ for all $a\in\Gamma$, which is equivalent to $az=za$ for all homogeneous $a\in\Gamma_{j}$. Decompose $z=\sum_{i\in\mathbb{Z}}z_{i}$ into its homogeneous components, and note that
$\sum_{i\in\mathbb{Z}}az_{i}=az=za=\sum_{i\in\mathbb{Z}}z_{i}a$
has $az_{i}$ as its homogeneous component of degree $i+j$. But this component is also $z_{i}a$, and hence $z_{i}$ is central. Therefore, the center is generated by homogeneous elements, and hence is a graded subalgebra. When we refer to a grading of the center of a graded algebra, we are referring to this grading if not specified otherwise.
To alleviate notation, we write monomials as $\mathbf{x}^{e}=\prod_{i=1}^{n+1}x_{i}^{e_{i}}\in R$, where $e\in\mathbb{N}^{n+1}$.
Proposition 4.3.
Let $C$ be a periodic cut of type $\gamma$, and $\mathbf{x}^{e}\in R^{G}$ be an invariant monomial. Then the degree of $\mathbf{x}^{e}$ with respect to the grading induced by $C$ is
$|\mathbf{x}^{e}|=\frac{\langle e,\gamma\rangle_{n+1}}{m}.$
In particular, two periodic cuts have the same type if and only if they induce the same grading on $R^{G}$.
Proof.
For every vertex $v\in Q_{0}$, denote by $c_{v}$ a cycle starting at $v$ and consisting of $e_{i}$ many arrows of type $i$. Then the element $\mathbf{x}^{e}$ corresponds via $R\ast G\simeq kQ/I$ to $\sum_{v\in Q_{0}}c_{v}$. We recall from 2.9 that each such cycle $c_{v}$ has the same degree, so it suffices to compute the degree for one such cycle. We fix without loss of generality the cycle $c=c_{0}$ which passes through 00. Then we view $c$ instead on $\hat{Q}$ as a path $p$ of length $l=\langle e,\mathbf{1}\rangle_{n+1}$ from 00 to $y$, where necessarily $y\in L_{1}$. It follows that
$(e_{1}-e_{n+1},e_{2}-e_{n+1},\ldots,e_{n}-e_{n+1},0)=e-e_{n+1}\mathbf{1}=(y,0).$
We note that $h_{C}(y)$ can be computed as
$h_{C}(y)=l-(n+1)\cdot|c|,$
since $|c|$ is the number of cut arrows in $p$. Using 2.9, we therefore have
$h_{C}(y)=\left\langle y,\mathbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}=l-(n+1)\cdot|c|.$
Then we rewrite
$\left\langle y,\mathbf{1}-\frac{n+1}{m}\gamma\right\rangle_{n}=\langle e-e_{n+1}\mathbf{1},\mathbf{1}\rangle_{n}-\frac{n+1}{m}\langle e-e_{n+1}\mathbf{1},\gamma\rangle_{n}.$
Thus, we have
$|c|$ $=\frac{1}{n+1}\left(\langle e,\mathbf{1}\rangle_{n+1}-\langle e-e_{n+1}\mathbf{1},\mathbf{1}\rangle_{n}+\frac{n+1}{m}\langle e-e_{n+1}\mathbf{1},\gamma\rangle_{n}\right)$
$=\frac{1}{n+1}\left(e_{n+1}+\langle e-e+e_{n+1}\mathbf{1},\mathbf{1}\rangle_{n}+\frac{n+1}{m}\langle e-e_{n+1}\mathbf{1},\gamma\rangle_{n}\right)$
$=\frac{1}{n+1}\left(e_{n+1}+ne_{n+1}+\frac{n+1}{m}(\langle e,\gamma\rangle_{n}-e_{n+1}\langle\mathbf{1},\gamma\rangle_{n})\right)$
$=e_{n+1}+\frac{1}{m}\left(\langle e,\gamma\rangle_{n}-e_{n+1}\langle\mathbf{1},\gamma\rangle_{n}\right)$
$=e_{n+1}+\frac{1}{m}\left(\langle e,\gamma\rangle_{n}-e_{n+1}(m-\gamma_{n+1})\right)$
$=\frac{1}{m}\langle e,\gamma\rangle_{n+1},$
which proves our claim. Note that the formula for the degree of $\mathbf{x}^{e}$ only depends on the type of the cut, and that $\gamma$ can be reconstructed from the degrees of the pure powers in $R^{G}$, so two cuts define the same grading if and only if they have the same type. ∎
The above proposition therefore means that we can identify the types of cuts with certain gradings on $R^{G}$. Since we are in the toric case, this has an especially nice interpretation.
Remark 4.4.
Note that $\mathbb{Z}$-gradings of a coordinate ring $k[X]$ correspond to $k^{\ast}$-actions on the affine variety $X=\operatorname{Spec}(R[X])$ by [SGA3, Proposition 4.7.3]. In detail, a $\mathbb{Z}$-grading on $k[X]$ is equivalent to a coaction morphism $k[X]\to k[X]\otimes_{k}k[t,t^{-1}]$ for the ring of Laurent polynomials. Taking the spectrum on both sides then gives an action of $k^{\ast}=\operatorname{Spec}(k[t,t^{-1}])$ on $X$. Since our variety $X=\operatorname{Spec}(R^{G})$ is toric, with a fixed torus $\mathbb{T}$, and because the generating monomials of $R^{G}$ are homogeneous for the gradings we consider, we obtain that these gradings on $R^{G}$ correspond to $1$-parameter subgroups of $(X,\mathbb{T})$. Furthermore, $X$ is defined by the cone $C\subseteq N_{\mathbb{R}}$, where $N=\operatorname{Hom}(k^{\ast},\mathbb{T})$ is the lattice of $1$-parameter subgroups. We note that this lattice is isomorphic to the lattice we considered in Section 3.2.
Proposition 4.5.
Let $P$ be the lattice simplex of types of cuts on $R\ast G\simeq kQ/I$, and $N$ the lattice of $1$-parameter subgroups of the toric variety $X=\operatorname{Spec}(R^{G})$. Then the lattice simplex in $N$ given by intersecting the defining cone $C\subseteq N_{\mathbb{R}}$ with the hyperplane corresponding to the degree condition $|\mathbf{x}^{\mathbf{1}}|=|\prod_{i=1}^{n+1}x_{i}|=1$, is isomorphic to $P$.
Proof.
Using 4.3, we identify types with the induced gradings, and with lattice points in $N$. Under this identification, we have for any type $\gamma$ that $|\mathbf{x}^{\mathbf{1}}|=\frac{\langle\mathbf{1},\gamma\rangle_{n+1}}{m}=1$, so the types lie on the hyperplane specified by this degree condition. To see that the lattice point $\lambda$ corresponding to a type lies in the cone $C$, it suffices to check that $\lambda$, seen as a $1$-parameter subgroup, converges. But this is immediate since the grading assigns non-negative degrees $d_{i}\geq 0$ to any invariant monomial, so in appropriately chosen coordinates, $\lambda$ acts by multiplication by $t^{d_{i}}$ for $t\in k^{\ast}$. ∎
With these gradings, the Gorenstein ring $R^{G}$ becomes graded so that its $a$-invariant is $a=-1$. This corresponds precisely to the condition in 1.4 that $R\ast G$ has Gorenstein parameter $1$.
Proposition 4.6.
Let $R^{G}$ be as before, graded by some cut. Then the graded canonical module of $R^{G}$ is the ideal generated by $(\Pi_{i=1}^{n+1}x_{i})$. In particular, as a graded Gorenstein ring, $R^{G}$ has $a$-invariant $a=-1$.
Proof.
The fact that $R^{G}$ is Gorenstein is well known, since $G\leq\operatorname{SL}_{n+1}(k)$ [Watanabe]. In particular, the graded canonical module $\omega$ is generated by a single element. By [BrunsHerzog, Theorem 6.3.5], $\omega$ is generated by the monomials in the interior of $C^{\vee}$, and in our case it is clear that this is generated by $(\Pi_{i=1}^{n+1}x_{i})$. By 4.3, we have that $|\Pi_{i=1}^{n+1}x_{i}|=1$ for any cut, and therefore $\omega(-1)\simeq R$. Hence $a=-1$. ∎
For completeness, we give another interpretation more directly in line with the McKay correspondence: The types of cuts correspond precisely to junior elements in $G$. This brings us back to the original construction from [AIR], since we made explicit use of the juniors in 3.1 and hence in 3.4 to construct our cuts.
Remark 4.7.
In [ItoReid, Section 2.2], a different parametrisation of the lattice of $1$-parameter subgroups is chosen. To do this, we recall that we have diagonalised the action of $G$, so that every element can be written as $g=\frac{1}{|g|}(e_{1},\ldots,e_{n+1})$. Of course, this notation depends on a choice of root of unity, but if we fix the diagonalisation and the root of unity, we can identify $N$ with the lattice $N_{G}$ generated by $\mathbb{Z}^{n+1}+\{\frac{1}{|g|}(e_{1},\ldots,e_{n+1})^{t}\mid g\in G\}$ in $\mathbb{Q}^{n+1}$, where we now view $\frac{1}{|g|}(e_{1},\ldots,e_{n+1})^{t}$ as a column vector with rational entries. An element $g=\frac{1}{|g|}(e_{1},\ldots,e_{n+1})$ is called junior if $\frac{1}{|g|}(e_{1}+\ldots+e_{n+1})=1$, and our simplex $P$ of types of cuts is then isomorphic to the lattice simplex of junior elements in $N_{G}$. Note that this means that the internal points of $P$ correspond to the junior elements of $G$.
We conclude by pointing out that by general toric geometry, the lattice vectors in $P$ respectively the vectors of junior elements in $N_{G}$ correspond to crepant divisors.
Remark 4.8.
Let $N_{G}$ be as above, and denote $C_{G}\subseteq(N_{G})_{\mathbb{R}}$ the cone defining $X=\operatorname{Spec}(R^{G})$. By [ItoReid], the lattice points in $C_{G}$ corresponding to junior elements in $G$ are in bijection with crepant exceptional prime divisors of any resolution of $X$. Furthermore, it was shown in [YamagishiIAbelian, Section 3.2, Section 6] that the rays generated by these lattice vectors in $C_{G}$ correspond to exceptional crepant divisors in a relative minimal model.
5. Mutation lattices
In this section, we recall mutation of cuts, show that cuts of a given type can be turned into a finite distributive lattice whose cover relations coincide with (nonzero) mutations, and hence that any two cuts of the same type are related by a sequence of mutations. We also give an explicit construction of the maximal element in the lattice. As before, we fix a cofinite $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$ and $m=|L_{0}/L_{1}|$.
5.1. Mutation
In this subsection, we provide a summary of cut mutation and its relevance to the $n$-representation infinite algebras we consider. This mutation is a higher version of the reflection functors of Berstein, Gel’fand and Ponomarev [BGP]. In modern language, these functors are rephrased via Auslander-Platzeck-Reiten tilting modules [APR], and we refer to $n$-APR tilting as introduced in [IyamaOppermann]. We fix a cut $C$ on $Q$, and recall that $Q_{C}=(Q_{0},Q_{1}-C)$.
Construction 5.1.
We call a vertex $x$ in $Q_{C}$ mutable if it is a source or a sink in $Q_{C}$. If $x$ is a source, that means that all arrows $a\in Q_{1}$ with $t(a)=x$ are in $C$. We note that all arrows $a\in Q_{1}$ with $s(a)=x$ are not in $C$, since otherwise we could complete such an $a\in C$ into an elementary cycle, whose last arrow is also in $C$ by assumption. Therefore, we can construct a new cut
$\mu_{x}(C)=\{a\in C\mid s(a)\neq x\}\cup\{a\in Q\mid t(a)=x\}$
which we call the mutation of $C$ at $x$. Mutation at a sink is defined dually. We further call this a nonzero mutation if $x\neq 0$. We also use the same terminology for the infinite cut quiver $\hat{Q}_{C}$. In particular, when mutating at a source $x$ in $\hat{Q}_{C}$, we mutate at all points $x+L_{1}$ simultaneously to preserve periodicity.
Clearly, mutation at a source $x$ in $Q_{C}$ turns it into a sink in $Q_{\mu_{x}(C)}$ and vice versa.
Remark 5.2.
Cut mutation as defined above is interesting because it corresponds to $n$-APR tilting of the $n$-representation infinite algebra $kQ_{C}/I_{C}$, as defined in [IyamaOppermann]. In particular, cut mutation from $C$ to $\mu_{x}(C)$ gives rise to a derived equivalence between the $n$-representation infinite algebras $kQ_{C}/I_{C}$ and $kQ_{\mu_{x}(C)}/I_{\mu_{x}(C)}$.
We also note that nonzero mutations have very nice effect on height functions.
Remark 5.3.
Let $C$ be a periodic cut on $\hat{Q}$, and $h_{C}$ the associated height function. If $0\neq s\in\hat{Q}_{C}$ is a source then
$h_{\mu_{s}(C)}(x)=\begin{cases}h_{C}(x)+n+1,&x-s\in L_{1}\\ h_{C}(x),&\text{else}.\end{cases}$
5.2. Lattices
The goal of this section is to show that for any type $\gamma$, the set $\{C\mid C\text{ is }L_{1}\text{-periodic and }\theta(C)=\gamma\}$ is a finite distributive lattice whose cover relations are given by nonzero mutations. This insight comes from the heuristic that we normalised height functions to $h_{C}(0)=0$, so we “should not” mutate at 00. As a corollary, we show that any two cuts of the same type are related by a sequence of mutations. As before, we fix a cofinite $L_{1}\xhookrightarrow{B^{\prime}}L_{0}$ and a type $\gamma$. Furthermore, we enumerate the elements of $L_{0}/L_{1}$ by $\{1,\ldots,m\}$ to fix an isomorphism $\mathbb{Z}^{|L_{0}/L_{1}|}\simeq\mathbb{Z}^{m}$. The relevant background on poset and lattice theory can be found in [Stanley, Chapter 3], and we use the standard notation established there. The following is well known.
Proposition 5.4.
$(\mathbb{Z}^{m},\leq)$ is a distributive lattice with $x\vee y=\max(x,y)$ and $x\wedge y=\min(x,y)$, where $\leq$ and $\max$ and $\min$ are taken componentwise.
We construct a function which generalises the relative height function of two perfect matchings in a dimer model.
Proposition 5.5.
Let $C_{1}$ and $C_{2}$ be two $L_{1}$-periodic cuts of type $\gamma$ and let $h_{C_{1}}$ and $h_{C_{2}}$ be the corresponding height functions. Then $\frac{h_{C_{1}}-h_{C_{2}}}{n+1}$ is an integer valued $L_{1}$-periodic function on $L_{0}$.
Proof.
It is clear that $(n+1)\mid(h_{C_{1}}(x)-h_{C_{2}}(x))$ for any $x\in L_{0}$ and any pair of cuts regardless of their types. Hence $\frac{h_{C_{1}}-h_{C_{2}}}{n+1}$ is integer valued. Let $x\in L_{0}$ and $y\in L_{1}$. Since $\theta(C_{1})=\theta(C_{2})=\gamma$, we know from 2.9 that $h_{C_{1}}(y)=h_{C_{2}}(y)=h_{\gamma}(y)$. Therefore:
$(h_{C_{1}}-h_{C_{2}})(x+y)$ $=h_{C_{1}}(x+y)-h_{C_{2}}(x+y)$
$=h_{C_{1}}(x)+h_{C_{1}}(y)-h_{C_{2}}(x)-h_{C_{2}}(y)$
$=h_{C_{1}}(x)+h_{\gamma}(y)-h_{C_{2}}(x)-h_{\gamma}(y)$
$=h_{C_{1}}(x)-h_{C_{2}}(x)=(h_{C_{1}}-h_{C_{2}})(x),$
which shows $L_{1}$-periodicity of $\frac{h_{C_{1}}-h_{C_{2}}}{n+1}$. ∎
Construction 5.6.
Let $C_{0}$ be an $L_{1}$-periodic cut of type $\gamma$. We will fix it and call it the reference cut. Then for any $L_{1}$-periodic cut $C$ of type $\gamma$, the function
$\overline{h}_{C}=\overline{h}_{C,C_{0}}\colon L_{0}/L_{1}\to\mathbb{Z},x+L_{1}\mapsto\frac{h_{C}(x)-h_{C_{0}}(x)}{n+1}$
is well-defined and called the relative height function of $C$ with respect to $C_{0}$. We suppress the dependence on $C_{0}$ whenever possible. Therefore, any $C$ can be identified with the corresponding vector $v_{C}=(\overline{h}_{C}(x+L_{1}))_{x+L_{1}\in L_{0}/L_{1}}\in\mathbb{Z}^{m}$. Clearly, the assignment $C\mapsto v_{C}\in\mathbb{Z}^{m}$ is injective and this way the set of all $L_{1}$-periodic cuts of type $\gamma$ becomes a subset of $\mathbb{Z}^{m}$.
We will now show that the set of vectors $v_{C}$ is closed under taking $\min$ and $\max$, in other words that this subset is a sublattice of $\mathbb{Z}^{m}$.
Proposition 5.7.
Let $C_{1}$ and $C_{2}$ be two $L_{1}$-periodic cuts of type $\gamma$ and let $h_{C_{1}}$ and $h_{C_{2}}$ be the corresponding height functions. Then the functions $h_{\min}=\min(h_{C_{1}},h_{C_{2}})$ and $h_{\max}=\max(h_{C_{1}},h_{C_{2}})$, are $L_{1}$-equivariant height functions. Moreover, $C_{h_{\min}}$ and $C_{h_{\max}}$ have type $\gamma$.
Proof.
We will only prove the claim for $h_{\min}$ since the proof for $h_{\max}$ is analogous. It is obvious that $h_{\min}(0)=0$. Secondly, we will show that $h_{\min}(x+\alpha_{i})\in\{h_{\min}(x)+1,h_{\min}(x)-n\}$ for any $x\in L_{0}$ and any $i\in\{1,\ldots,n+1\}$. Fix such an $x$ and such an $i$. We will assume without loss of generality that $h_{C_{1}}(x)\leq h_{C_{2}}(x)$. If $h_{C_{1}}(x)=h_{C_{2}}(x)$, we are done since $h_{\min}$, restricted to points $x$ and $x+\alpha_{i}$, equals either $h_{C_{1}}$ or $h_{C_{2}}$, restricted to the same pair of points. Therefore we will further assume $h_{C_{1}}(x)<h_{C_{2}}(x)$. If $h_{C_{1}}(x+\alpha_{i})\leq h_{C_{2}}(x+\alpha_{i})$, we are done again by the same argument as before. Therefore, the only interesting case is $h_{C_{1}}(x)<h_{C_{2}}(x)$ and $h_{C_{1}}(x+\alpha_{i})>h_{C_{2}}(x+\alpha_{i})$. These two conditions imply $h_{C_{1}}(x+\alpha_{i})=h_{C_{1}}(x)+1$ and $h_{C_{2}}(x+\alpha_{i})=h_{C_{2}}(x)-n$. But then $h_{C_{1}}(x)<h_{C_{2}}(x)=h_{C_{2}}(x+\alpha_{i})+n<h_{C_{1}}(x+\alpha_{i})+n=h_{C_{1}}(x)+1+n$, which in particular gives $h_{C_{1}}(x)<h_{C_{2}}(x)<h_{C_{1}}(x)+1+n$, in other words, $10<h_{C_{2}}(x)-h_{C_{1}}(x)<n+$ 1. This is a contradiction since for any pair of cuts $C_{1}$ and $C_{2}$ and for any point $x\in L_{0}$ we have $(n+1)|(h_{C_{2}}(x)-h_{C_{1}}(x))$.
Next, we need to show equivariance of $h_{\min}$. Let $x\in L_{0}$ and $y\in L_{1}$. Since $\theta(C_{1})=\theta(C_{2})=\gamma$, we know from 2.9 that $h_{C_{1}}(y)=h_{C_{2}}(y)=h_{\gamma}(y)$. Therefore,
$h_{\min}(x+y)$ $=\min(h_{C_{1}}(x+y),h_{C_{2}}(x+y))$
$=\min(h_{C_{1}}(x)+h_{\gamma}(y),h_{C_{2}}(x)+h_{\gamma}(y))$
$=\min(h_{C_{1}}(x),h_{C_{2}}(x))+h_{\gamma}(y)$
$=h_{\min}(x)+h_{\min}(y).$
Finally, the height function $h_{\min}$ corresponds to a cut of type $\gamma$, which follows from 2.12. ∎
The above clearly translates to relative height functions and hence to the vectors $v_{C}$. Note that there are finitely many $L_{1}$-periodic cuts of a given type, hence we obtain the following.
Corollary 5.8.
The set $\{v_{C}\mid C\text{ is an }L_{1}\text{-periodic cut, }\theta(C)=\gamma\}$ is a finite sublattice of $(\mathbb{Z}^{m},\leq)$. Therefore, it is a finite distributive lattice.
We transfer notation along the bijection $C\mapsto v_{C}$.
Remark 5.9.
It follows that the set $\{C\mid C\text{ is an }L_{1}\text{-periodic cut, }\theta(C)=\gamma\}$ is a finite distributive lattice, where we write $C_{1}\preceq C_{2}$ iff $v_{C_{1}}\leq v_{C_{2}}$. The join and meet of two cuts $C_{1}$ and $C_{2}$ is given by the cuts $C_{1}\vee C_{2}$ and $C_{1}\wedge C_{2}$ defined by the height functions $h_{\max}$ and $h_{\min}$ from 5.7. These satisfy $v_{C_{1}\wedge C_{2}}=\min(v_{C_{1}},v_{C_{2}})$ and $v_{C_{1}\vee C_{2}}=\max(v_{C_{1}},v_{C_{2}})$. We therefore refer to the lattice of cuts of type $\gamma$ from now on.
Next, we will explore the cover relations of the lattice in question. Recall that a cover relation is a relation of the form $C_{1}\prec C_{2}$ such that there is no $C_{3}$ with $C_{1}\prec C_{3}\prec C_{2}$. A cover relation is denoted by $C_{1}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$. From now on we will mostly be interested in the case when $\gamma$ is positive, in other words, $\gamma_{i}>0$ for all $i\in\{1,\ldots,n+1\}$. In this case cover relations have a particularly nice interpretation. We will first briefly note which effect a mutation of a cut $C$ has on the vector $v_{C}$.
Remark 5.10.
Let $C_{1}$ and $C_{2}$ be two $L_{1}$-periodic cuts of a positive type $\gamma$. Assume $C_{1}$ can be obtained from $C_{2}$ via mutation at a nonzero sink $s$. Then $v_{C_{1}}=v_{C_{2}}-e_{i}$, where $e_{i}$ is the $i$th standard vector, with $i$ being the position indexing the point $s+L_{1}$.
Theorem 5.11.
Let $C_{1}$ and $C_{2}$ be two $L_{1}$-periodic cuts of a positive type $\gamma$. Then $C_{1}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$ if and only if $C_{1}$ can be obtained from $C_{2}$ via a nonzero sink mutation (or, equivalently, $C_{2}$ can be obtained from $C_{1}$ via a nonzero source mutation).
Proof.
First, we assume that $C_{1}$ can be obtained from $C_{2}$ by a mutation at a nonzero sink $s$. Then by 5.10 we have $v_{C_{1}}=v_{C_{2}}-e_{i}$ for some $i$. Since there are no integer vectors strictly between $v_{C_{1}}$ and $v_{C_{2}}$, it follows that $C_{1}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$. Conversely, assume that $C_{1}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$, then in particular it follows that $h_{C_{1}}(x)\leq h_{C_{2}}(x)$ for all $x\in L_{0}$. If we have $h_{C_{1}}(x)=h_{C_{2}}(x)$ for all $x\in L_{0}$ then $C_{1}=C_{2}$, which is a contradiction, hence we fix some $x\in L_{0}$ such that $h_{C_{1}}(x)<h_{C_{2}}(x)$. Next, recall that it follows from 2.14 that there are no infinite paths in the infinite cut quiver $\hat{Q}_{C_{2}}$. We follow any path starting at $x$ in $\hat{Q}_{C_{2}}$, and this path must be finite, which means it ends at a sink $s\in L_{0}$. Since we followed a path in $\hat{Q}_{C_{2}}$, the height function $h_{C_{2}}$ increased along each arrow, so we have $h_{C_{1}}(s)<h_{C_{2}}(s)$. Note that $s\not\in L_{1}$, since otherwise 2.9 gives $h_{C_{1}}(s)=h_{\gamma}(s)=h_{C_{2}}(s)$. That means we can perform a nonzero sink mutation of $C_{2}$ at $s$ and obtain a new cut $C_{3}$. Note that $h_{C_{1}}(x)\leq h_{C_{2}}(x)$ for all $x\in L_{0}$, combined with $h_{C_{1}}(s)<h_{C_{2}}(s)$, implies $v_{C_{1}}+e_{i}\leq v_{C_{2}}$. Here $e_{i}$ is the $i$th standard vector with $i$ being the position indexing point $s+L_{1}$. Together with that, 5.10 gives $v_{C_{3}}=v_{C_{2}}-e_{i}$ with the same index $i$, hence we get $v_{C_{1}}\leq v_{C_{3}}<v_{C_{2}}$. Since we assumed $C_{1}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$, we cannot have $v_{C_{1}}<v_{C_{3}}<v_{C_{2}}$, hence $v_{C_{1}}=v_{C_{3}}$. It follows that $C_{1}=C_{3}$, which by construction was obtained from $C_{2}$ by a single nonzero sink mutation. ∎
Corollary 5.12.
Any pair of cuts $C_{1}$ and $C_{2}$ of a positive type $\gamma$ can be obtained one from another via a sequence of (nonzero) mutations.
Proof.
Since $C_{1}\wedge C_{2}\preceq C_{1}$ and $C_{1}\wedge C_{2}\preceq C_{2}$, there are chains of covering relations $C_{1}\mathrel{\cdot}\mathrel{\mkern-3.0mu}\succ\ldots\mathrel{\cdot}\mathrel{\mkern-3.0mu}\succ C_{1}\wedge C_{2}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}\ldots\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}C_{2}$. Each of these relations corresponds to a nonzero mutation and thus the statement holds. ∎
Let us note an interesting consequence of the above.
Remark 5.13.
In [AIR], Amiot, Iyama and Reiten construct a cut for the skew-group algebra $R\ast G$ where $G$ is a cyclic group generated by a junior element. This is the same construction we used in 3.1. They furthermore note that it would be interesting to generalise their result to non-cyclic groups. Our results show that for abelian groups, there are no other cuts up to $n$-APR tilting and skewing. More precisely, it follows from 3.4 that the cuts we first construct arise by decomposing the abelian group $G\simeq H\times K$ so that $H$ is cyclic. Then $R\ast H$ is given a cut via Amiot-Iyama-Reiten’s procedure, and this is extended to $R\ast G$ via the isomorphism $(R\ast G)\simeq(R\ast H)\ast K$, see [DramburgGasanova, Section 5] for details. Then 5.12 shows that every cut is mutation equivalent to one induced from Amiot-Iyama-Reiten’s procedure, and hence any $n$-representation infinite algebra $\Lambda$ is $n$-APR tilting equivalent to some $\Lambda^{\prime}\ast K$ where $K$ is an abelian group acting on the $n$-representation infinite algebra $\Lambda^{\prime}$ obtained from Amiot-Iyama-Reiten’s construction.
5.3. Constructing the minimal and the maximal element of the mutation lattice for a positive type
For a fixed type $\gamma$, the mutation lattice is finite and distributive. Therefore, it has a unique maximal element, the construction of which is the goal of this subsection. In a dual procedure, one can construct the unique minimal element. We fix a type $\gamma$, and use the following notation throughout this subsection: for $x\in\mathbb{Z}^{n}$ and $z\in\mathbb{Z}$ we will write $(x,0)=(x_{1},\ldots,x_{n},0)$ and $u_{z}(x)=(x,0)-z\textbf{1}$. Further, we consider the linear map $\psi(y)=(y,0)-\frac{\langle y,\gamma\rangle_{n}}{m}\mathbf{1}$ and write $L=\psi(L_{1})\subset\mathbb{Z}^{n+1}$. Note that $\langle l,\gamma\rangle_{n+1}=0$ for any $l\in L$. For every $x\in L_{0}$, let $S_{x}=\{z\in\mathbb{Z}\mid\exists l\in L:u_{z}(x)\geq l\}$.
Construction 5.14.
For every $x\in L_{0}$, the set $S_{x}$ is nonempty and bounded above. Therefore, we can construct the function
$p\colon L_{0}\to\mathbb{Z},x\to\max S_{x}.$
Proof.
To see that $S_{x}$ is nonempty, simply note that by taking $z$ small enough, one can always reach a situation when $u_{z}(x)\geq(0,\ldots,0)\in L$. Now we show that $S_{x}$ is bounded above. Indeed, there exists $z_{0}$ such that for any $z\geq z_{0}$ all the coordinates of $u_{z}(x)$ are negative. Fix any $z_{1}\geq z_{0}$. We claim that $z_{1}\not\in S_{x}$. Assume on the contrary that $u_{z_{1}}(x)\geq l$ for some $l\in L$. Then $u_{z_{1}}(x)=l+\varepsilon$, where $\varepsilon$ is some vector with nonnegative coordinates. Then $0=\langle l,\gamma\rangle_{n+1}=\langle u_{z_{1}}(x)-\varepsilon,\gamma\rangle_{n+1}$, which is a contradiction since all the coordinates of $u_{z_{1}}(x)-\varepsilon$ are negative, and $\gamma$ is a nonzero nonnegative vector. Therefore, $S_{x}$ is a nonempty bounded above set of integers. Therefore, it has a maximum and hence the function $p$ is well defined. ∎
Next, we need to show the following properties of $p$.
Lemma 5.15.
For $y\in L_{0}$, we have $(1)\Rightarrow(2)\Rightarrow(3)$, where
  1. (1)
    $u_{p(y)}(y)\in L$,
  2. (2)
    $y\in L_{1}$,
  3. (3)
    $p(y)=\frac{\langle y,\gamma\rangle_{n}}{m}$.
Proof.
$(1)\Rightarrow(2)$: Let $u_{p(y)}(y)\in L$. Then $u_{p(y)}(y)=\psi(y^{\prime})$ for some $y^{\prime}\in L_{1}$, which gives
$(y,0)-p(y)\textbf{1}=(y^{\prime},0)-\frac{\langle y^{\prime},\gamma\rangle_{n}}{m}\textbf{1},$
which, if one compares the last coordinates, gives $p(y)=\frac{\langle y^{\prime},\gamma\rangle_{n}}{m}$ and hence $(y,0)=(y^{\prime},0)$ and thus $y=y^{\prime}\in L_{1}$.
$(2)\Rightarrow(3)$: Let $y\in L_{1}$. Then
$\hskip 22.76228ptp(y)$ $=\max\{z\in\mathbb{Z}\mid\exists l\in L:(y,0)-z\textbf{1}\geq l\}$
$=\max\left\{z\in\mathbb{Z}\mid\exists l\in L:(y,0)-\left(z+\frac{\langle y,\gamma\rangle_{n}}{m}\right)\textbf{1}\geq l\right\}+\frac{\langle y,\gamma\rangle_{n}}{m}$
$=\max\left\{z\in\mathbb{Z}\mid\exists l\in L:\psi(y)-z\textbf{1}\geq l\right\}+\frac{\langle y,\gamma\rangle_{n}}{m}$
$=\max\left\{z\in\mathbb{Z}\mid\exists l^{\prime}\in L-\psi(y):-z\textbf{1}\geq l^{\prime}\right\}+\frac{\langle y,\gamma\rangle_{n}}{m}$
$=\max\left\{z\in\mathbb{Z}\mid\exists l\in L:-z\textbf{1}\geq l\right\}+\frac{\langle y,\gamma\rangle_{n}}{m}.$
Now, we claim that $\max\left\{z\in\mathbb{Z}\mid\exists l\in L:-z\textbf{1}\geq l\right\}=0$. Indeed, if $z_{0}\in\mathbb{Z}$ belongs to this set, we get $-z_{0}\textbf{1}=l+\varepsilon$, where $\varepsilon$ is a nonnegative vector, and then
$-z_{0}m=\langle-z_{0}\textbf{1},\gamma\rangle_{n+1}=\langle l,\gamma\rangle_{n+1}+\langle\varepsilon,\gamma\rangle_{n+1}=\langle\varepsilon,\gamma\rangle_{n+1}\geq 0,$
which gives $z_{0}\leq 0$. Clearly, 00 belongs to the above set, and thus we get $p(y)=\frac{\langle y,\gamma\rangle_{n}}{m}$. ∎
Remark 5.16.
In fact, it is possible to show that $(1)\Leftrightarrow(2)\Rightarrow(3)$ in 5.15, and that the remaining implication $(3)\Rightarrow(1)$ holds if additionally $\gamma$ is assumed to be positive. However, we only ever need the implications $(1)\Rightarrow(2)\Rightarrow(3)$.
Now we are ready to define a candidate height function for our maximal element.
Theorem 5.17.
Let $\gamma$ be a (not necessarily positive) type and let
$h\colon L_{0}\to\mathbb{Z},x\mapsto\langle u_{p(x)}(x),\mathbf{1}\rangle_{n+1}.$
Then $h$ is an $L_{1}$-equivariant height function and $C_{h}$ is of type $\gamma$.
Proof.
We first show that $h$ is a height function. Since $0\in L_{1}$, the implication $(2)\Rightarrow(3)$ of 5.15 tells us that $p(0)=0$ and thus $h(0)=0$. Now we want to compare $h(x+\alpha_{i})$ with $h(x)$ for all $i\in\{1,\ldots,n+1\}$. First we will consider the case $i\in\{1,\ldots,n\}$, where we can without loss of generality assume $i=1$. Let $i=1$. We will first show that $p(x+\alpha_{1})\in\{p(x),p(x)+1\}$. Indeed,
$p(x)$ $=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}\geq l\}$
$\leq\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}+(1,0,\ldots,0)\geq l\}$
$\leq\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}+\textbf{1}\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-(z-1)\textbf{1}\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}\geq l\}+1=p(x)+1.$
Since $p(x+\alpha_{1})=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}+(1,0,\ldots,0)\geq l\}$, we conclude that $p(x+\alpha_{1})\in\{p(x),p(x)+1\}$. Recall that
$h(x+\alpha_{1})-h(x)=\langle u_{p(x+\alpha_{1})}(x+\alpha_{1})-u_{p(x)}(x),\textbf{1}\rangle_{n+1}.$
If $p(x+\alpha_{1})=p(x)$, we get
$u_{p(x+\alpha_{1})}(x+\alpha_{1})-u_{p(x)}(x)$ $=u_{p(x)}(x+\alpha_{1})-u_{p(x)}(x)$
$=u_{p(x)}(x)+(1,0,\ldots,0)-u_{p(x)}(x)=(1,0,\ldots,0)$
and hence $h(x+\alpha_{1})-h(x)=1$. If $p(x+\alpha_{1})=p(x)+1$, we get
$u_{p(x+\alpha_{1})}(x+\alpha_{1})-u_{p(x)}(x)$ $=u_{p(x)+1}(x+\alpha_{1})-u_{p(x)}(x)$
$=u_{p(x)+1}(x)+(1,0,\ldots,0)-u_{p(x)}(x)$
$=u_{p(x)}(x)-\textbf{1}+(1,0,\ldots,0)-u_{p(x)}(x)=(0,-1,\ldots,-1)$
and hence $h(x+\alpha_{1})-h(x)=-n$.
The case $i=n+1$ can be treated analogously, showing that
$p\left(x-\left(\sum_{i=1}^{n}\alpha_{i}\right)\right)=p(x+\alpha_{n+1})\in\{p(x)-1,p(x)\},$
and that $h(x+\alpha_{n+1})-h(x)=1$ when $p(x+\alpha_{n+1})=p(x)-1$ and $h(x+\alpha_{n+1})-h(x)=-n$ when $p(x+\alpha_{n+1})=p(x)$. Therefore, we conclude that $h$ is a height function. Now we want to show that $h$ is $L_{1}$-equivariant. In order to show this, it is enough to show $L_{1}$-equivariance of $p(x)$. Indeed, $p(x+y)=p(x)+p(y)$ implies
$h(x+y)$ $=\langle u_{p(x+y)}(x+y),\textbf{1}\rangle_{n+1}$
$=\langle u_{p(x)+p(y)}(x+y),\textbf{1}\rangle_{n+1}=\langle u_{p(x)}(x)+u_{p(y)}(y),\textbf{1}\rangle_{n+1}$
$=\langle u_{p(x)}(x),\textbf{1}\rangle_{n+1}+\langle u_{p(y)}(y),\textbf{1}\rangle_{n+1}=h(x)+h(y).$
Now, the implication $(2)\Rightarrow(3)$ of 5.15 tells us that for any $y\in L_{1}$ we have $p(y)=\frac{\langle y,\gamma\rangle_{n}}{m}$ and hence $u_{p(y)}(y)=\psi(y)$. Therefore,
$p(x+y)$ $=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x+y,0)-z\textbf{1}\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-(z-p(y))\textbf{1}+(y,0)-p(y)\textbf{1}\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-(z-p(y))\textbf{1}+\psi(y)\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l^{\prime}\in L-\psi(y):(x,0)-(z-p(y))\textbf{1}\geq l^{\prime}\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-(z-p(y))\textbf{1}\geq l\}$
$=\max\{z\in\mathbb{Z}\mid\exists l\in L:(x,0)-z\textbf{1}\geq l\}+p(y)=p(x)+p(y).$
We conclude that $h(x)$ is an equivariant height function. Therefore, it defines an $L_{1}$-periodic cut $C_{h}$. As before, 5.15 gives us $u_{p(y)}(y)=\psi(y)$ for every $y\in L_{1}$, and we compute
$h(y)$ $=\langle u_{p(y)}(y),\textbf{1}\rangle_{n+1}=\left\langle(y,0)-\frac{\langle y,\gamma\rangle_{n}}{m}\textbf{1},\textbf{1}\right\rangle_{n+1}$
$=\langle(y,0),\textbf{1}\rangle_{n+1}-\frac{\langle y,\gamma\rangle_{n}}{m}\langle\textbf{1},\textbf{1}\rangle_{n+1}$
$=\langle y,\textbf{1}\rangle_{n}-\frac{n+1}{m}\langle y,\gamma\rangle_{n}=\left\langle y,\textbf{1}-\frac{n+1}{m}\gamma\right\rangle,$
hence $C_{h}$ has type $\gamma$ by 2.12. ∎
Proposition 5.18.
Let $h$ be as in 5.17. If $x$ is a source of $\hat{Q}_{C_{h}}$, then $x\in L_{1}$.
Proof.
Assume $x=(x_{1},\ldots,x_{n})$ is a source. Then all arrows $x-\alpha_{i}\to x$ for $i\in\{1,\ldots,n+1\}$ belong to $\hat{C}_{h}$. In other words, $h(x)-h(x-\alpha_{i})=-n$ for all $i\in\{1,\ldots,n+1\}$. From the proof of 5.17 we know that this happens if and only if $p(x)=p(x-\alpha_{i})+1$ for $i\in\{1,\ldots,n\}$ and $p(x)=p(x-\alpha_{n+1})$. By definition of $p(x)$ we know that there is some $l=(l_{1},\ldots,l_{n+1})\in L$ such that $(x,0)-p(x)\textbf{1}\geq l$. Since $p(x-\alpha_{1})=p(x)-1$, we get that $(x-\alpha_{1},0)-p(x)\textbf{1}$ can not be greater or equal to any element in $L$, in particular, we get
$l\not\leq(x-\alpha_{1},0)-p(x)\textbf{1}=(x,0)-p(x)\textbf{1}-(1,0,\ldots,0).$
Combining this with $l\leq(x,0)-p(x)\textbf{1}$ gives $l_{1}=x_{1}-p(x)$. The same holds for all $i\in\{1,\ldots,n\}$. Finally, we have $p(x)=p(x-\alpha_{n+1})$, which means $(x-\alpha_{n+1},0)-(p(x)+1)\textbf{1}$ can not be greater or equal to any element in $L$, in particular, we get
$l\not\leq(x+\alpha_{1}+\ldots+\alpha_{n},0)-(p(x)+1)\textbf{1}=(x,0)-p(x)\textbf{1}-(0,\ldots,0,1).$
Combining this with $l\leq(x,0)-p(x)\textbf{1}$ gives $l_{n+1}=-p(x)$. Therefore, we conclude that $u_{p(x)}(x)=(x,0)-p(x)\textbf{1}=l\in L$, and then the implication $(1)\Rightarrow(2)$ of 5.15 implies that $x\in L_{1}$. ∎
To show that $C_{h}$ is maximal, we need to assume that $\gamma$ is positive, since then mutations and cover relations coincide by 5.11.
Corollary 5.19.
If $\gamma$ is positive, then $C_{h}$ constructed in 5.17 is the unique maximal element in its mutation lattice. Furthermore, $Q_{C_{h}}$ has a unique source, located at 00.
Proof.
Assume that $C_{h}$ is not maximal, then there exists $D\succ C_{h}$ and a chain of cover relations $C_{h}\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}\ldots\prec\mathrel{\mkern-3.0mu}\mathrel{\cdot}D$. Then by 5.11, $Q_{C_{h}}$ has to have a nonzero source, which is a contradiction with 5.17. Finally, note that $Q_{C_{h}}$ is acyclic by 2.14, and since $Q_{C_{h}}$ is in addition finite, it must have a source, which must be located at 00. ∎
On the other hand, if $\gamma$ is not positive, the corresponding finite distributive lattice is harder to understand. Cover relations are not given by mutations, since none of the cut quivers have sources or sinks. The construction from 5.17 still produces a cut, but we are aware neither of a proof that it is the maximal element, nor of a counterexample.
Question 5.20.
Let $\gamma$ be a nonpositive type and let $C_{h}$ be the cut constructed in 5.17. Is it true that $C_{h}$ is the maximal element in the corresponding mutation lattice?
Acknowledgement
We thank Martin Herschend and David Ploog for many helpful discussions. The second author was supported by a fellowship from the Wenner-Gren Foundations (grant WGF2022-0052).
[…]